### v1

# Federated Learning Clients Clustering with Adaptation to Data Drifts

Abstract.
Federated Learning (FL) enables deep learning model training across edge devices and protects user privacy by retaining raw data locally. Data heterogeneity in client distributions slows model convergence and leads to plateauing with reduced precision. Clustered FL solutions address this by grouping clients with statistically similar data and training models for each cluster. However, maintaining consistent client similarity within each group becomes challenging when data drifts occur, significantly impacting model accuracy. In this paper, we introduce Fielding, a clustered FL framework that handles data drifts promptly with low overheads. Fielding detects drifts on all clients and performs selective label distribution-based re-clustering to balance cluster optimality and model performance, remaining robust to malicious clients and varied heterogeneity degrees. Our evaluations show that Fielding improves model final accuracy by 1.9%-5.9% and reaches target accuracies 1.16$\times$-2.61$\times$ faster.
copyright: nonedoi: isbn: price:
1. Introduction
Federated Learning (FL) is a distributed machine learning paradigm that enables multiple clients to train one model collaboratively while retaining raw data locally (Bonawitz et al., 2019; McMahan et al., 2017). As FL circumvents the communication costs of data movement and mitigates privacy leakage (Lyu et al., 2024; Huba et al., 2022), it has drawn attention across various fields and seen adoption in applications such as Google Keyboard (Yang et al., 2018), energy consumption forecasting (Abdulla et al., 2024), and medical image processing (Zhou et al., 2023).
The main challenge in federated learning is data heterogeneity. Since federated learning involves many devices of diverse settings, there exists substantial dataset size and distribution divergences between clients from distant geographical locations and with distinct personal habits (Li et al., 2022a; Kim et al., 2024). Such data heterogeneity leads to slow model convergence and plateauing with degraded precision (Li et al., 2020a; Zhao et al., 2018; Sattler et al., 2020). A wide range of clustering mechanisms have been proposed to group clients with similar data distribution into more homogeneous clusters and train one model for each cluster (Ghosh et al., 2020; Duan et al., 2021; Liu et al., 2023; Jothimurugesan et al., 2023).
Data drift occurs frequently in federated learning systems. It is defined as temporal changes in the label distribution (i.e., the proportion of samples assigned to each label) of clients’ local datasets. Such drifts arise from varying sampling rates, hardware configurations, sensor modalities, and limitations in client storage capacities (Chen et al., 2020; Nandi and Xhafa, 2022; Marfoq et al., 2023). Existing clustering mechanisms, however, are susceptible to data drift. Prior works address this issue but often either inadequately adjust clusters or incur prohibitive costs. IFCA and FlexCFL (Ghosh et al., 2020; Duan et al., 2021) adjust clusters by relocating shifted clients while maintaining a fixed number of clusters, operating under the assumption of low drift magnitude. Auxo (Liu et al., 2023) re-clusters clients participating in the current training round using gradient information; however, since only a subset of clients is selected each round, it delays adjustments for drifted clients until they are chosen for training. FedDrift (Jothimurugesan et al., 2023) re-clusters all clients and adaptively determines the number of clusters by transmitting all cluster models to every client each round to obtain training loss on local data, leading to significant communication and computation costs.
In this work, we propose Fielding, a novel clustering-based FL framework that retains cluster client similarity despite frequent data drifts. Fielding focuses on re-clustering clients as data drifts, dynamically adjusting the number of clusters while stabilizing responses to minor data drifts. Fielding combines per-client migration with selective global clustering to capture significant distribution patterns without compromising accuracy.
We assume the existence of a coordinator, typically deployed on a powerful central server to manage synchronization and advance global model training (Bonawitz et al., 2019; Bagdasaryan et al., 2020). Our solution handles all drifted clients by clustering them according to their local distribution vector, a piece of information readily available all the time. We present a theoretical framework supporting this approach. Fielding is designed to be lightweight; it introduces no additional computation on the client side and requires clients only to transmit their distribution vector (an array with up to a few thousand values) when they register with the coordinator.
We implement Fielding by extending the widely adopted FL engine FedScale (Lai et al., 2022) to support streaming data and our clustering mechanisms. We evaluated Fielding with up to 5078 clients on four image streaming traces with data drifts. Compared to existing approaches, Fielding improves model accuracy by 1.9% to 5.9% and achieves the target accuracy 1.16$\times$ to 2.61$\times$ faster. We also show that Fielding is compatible with various client selection and aggregation algorithms, and robust to malicious clients and varied heterogeneity degrees.
Our main contributions are:
  • We introduce Fielding, a clustered FL framework that handles data drifts promptly with low overheads.
  • We show that Fielding imposes minimal overhead by eliminating additional client-side computations and reducing communication costs through the use of readily available data, and we provide a theoretical framework to support our approach.
  • We extend the FedScale FL engine to implement Fielding and demonstrate its efficacy on four real-world datasets with up to 5,078 clients, achieving up to 5.9% accuracy improvement and faster convergence compared to existing methods.
  • We show that Fielding is compatible with various client selection and aggregation algorithms and is robust against malicious clients and different degrees of data heterogeneity.
2. Background
In this section, we provide the background of clustered federated learning and the impacts of data drifts.
2.1. Clustered Federated Learning
FL deployment often involves hundreds to thousands of available clients participating in the training (Bonawitz et al., 2019). For example, Google sets 100 as the target number of clients for a round of training when improving keyboard search suggestions (Yang et al., 2018) and (Huang et al., 2019) have developed mortality and stay time predictor for the eICU collaborative research database (Pollard et al., 2018) containing data of 208 hospitals and more than 200,000 patients. Although thousands of devices may be available in a given round, FL systems typically select a subset to participate in training. This selection process ensures a reasonable training time and accounts for diminishing returns where including more clients doesn’t accelerate convergence (Xu et al., 2024).
Data heterogeneity is a major challenge in FL. Varying volumes and distribution of local data among participants is known to cause slow convergence and poor model accuracy (Yang et al., 2021; Wen et al., 2023). To mitigate the impact of heterogeneity, an effective approach is clustering, which groups statistically similar clients. Clients in the same cluster train one cluster model collaboratively and also have their inference requests served by this model. Previous works such as Auxo (Liu et al., 2023), FlexCFL (Duan et al., 2021), FedDrift (Jothimurugesan et al., 2023), and FedAC (Zhang et al., 2024) have shown that clustering can achieve high accuracy and fast model convergence.
2.2. Data Drifts Affect Clustering
The key challenges for clustered federated learning are data drifts. Data drifts refer to the changes in the percentage of clients’ local samples associated with a specific label. Drifts happen naturally when clients have access to non-stationary streaming data (e.g., cameras on vehicles and virtual keyboards on phones) (Koh et al., 2021; Bhardwaj et al., 2022). For example, when training an FL model for keyboard suggestions, drifts occur when many clients simultaneously start searching for a recent event (widespread drift), a few clients pick up a niche hobby (concentrated intense drift), or a new term appears when a product is launched (new label added).
Data drifts make clustered FL ineffective as it increases the data heterogeneity in the clusters. We demonstrate this problem with the Functional Map of the World (FMoW) dataset (Christie et al., 2018). This dataset contains time-stamped satellite images labeled according to the function (e.g., recreational facility, crop field, etc.) of the captured land. We keep images taken on or after January 1, 2015, and create one client per unique UTM zone metadata value (302 clients in total). A day’s worth of data becomes available every two training rounds, and clients maintain images they received over the last 100 rounds.
We use mean client distance as the intra-cluster heterogeneity metric (Lai et al., 2021). For each client, we calculate the average pairwise L1 distance between its data distribution vector and those of all other clients in the same cluster. We then use the mean of these per-client distances to quantify the overall heterogeneity after clustering. Note that instead of first averaging within each cluster and then finding the mean of those cluster averages, we use the overall mean across all clients. This helps us avoid biased results arising from imbalanced cluster sizes. Note that in the baseline case without any clustering, we consider all clients as members of a “global cluster”.
Figure 1. Client heterogeneity of the global set and clusters generated through label-based clustering.
We cluster clients based on their label distribution vectors at the starting round and track the overall heterogeneity by measuring the mean client distance as described earlier. We cluster based on label distribution because data drifts occurs when label distribution changes. Figure 1 shows that at the beginning (e.g., round 1), clustering reduces the per-cluster heterogeneity compared with no clustering (i.e., putting all clients in a global set). However, as data distribution shifts over time, static clustering increases per-cluster heterogeneity and soon gets close to the no clustering case at round 322.
3. Fielding Design decisions
Addressing data drifts is crucial for preserving the similarity of clients within a cluster. To handle data drifts, we must identify which clients should be moved, determine the appropriate clusters for these clients, and specify the information to collect for drift detection and client movements.
3.1. Combine Per-client Adjustment and Re-clustering
Existing clustered FL works rely on per-client adjustment or re-clustering to handle data drifts. However, neither approach works well on all levels of drifts.
Per-client adjustment fails in many-drift scenarios. Per-client adjustment keeps a fixed number of clusters and moves drifted clients individually to another cluster with similar gradients or model weights (Ghosh et al., 2020; Duan et al., 2021). However, in highly dynamic environments where many clients change their distributions at the same time (Jothimurugesan et al., 2023), we have to move all these clients across clusters and the cluster membership changes dramatically. It is hard to decide which cluster the client should move to because the basis used by previous works to match client to the closest cluster (i.e., cluster’s average model weights or gradients) is changing (Ghosh et al., 2020; Duan et al., 2021). We demonstrate this problem in Figure 1 (“Move Individuals”). In rounds 1-40 and rounds 410-615, individual movements help keep the mean client distance below that of the global set even though clients are becoming more heterogeneous (signaled by the increasing heterogeneity value). However, as the global set of all clients experiences major drifts during rounds 600-730 (signaled by the heterogeneity spike), many clients start moving across clusters simultaneously. Per-client adjustment then results in poor clusters and an overall heterogeneity value higher than the global set in rounds 740-1340.
Moreover, maintaining a fixed number of clusters can lead to sub-optimal clustering when new distribution patterns emerge during data drifts. For instance, we want to create another cluster when a new label appears for clients with a substantial number of samples of that label, but keeping the cluster number static will deem such creation infeasible.
(a) Selective global re-clustering vs. default global re-clustering. Positive numbers mean selective re-clustering is better.
(b) Global re-clustering vs. Selected re-clustering. Positive numbers mean re-clustering all drifted clients is better.
Figure 2. Accuracy difference between different re-clustering approaches.
Re-clustering makes model unstable during small drifts In Figure 1, we show an ideal case where we get optimal clusters by globally re-clustering all clients using label distribution vectors when any client drifts (“Global recluster”). Our implementation bounds the maximum number of clusters to 10, as previous works empirically show that having more clusters can negatively affect model convergence due to limited training resources (Liu et al., 2023). As expected, this approach yields the lowest heterogeneity value. However, we will next show that such low heterogeneity does not necessarily lead to superior model accuracy.
Common clustering algorithms such as k-means have randomized initialization, leading to distinct results in multiple runs (Arthur and Vassilvitskii, 2006; Ahmed et al., 2020). This behavior is also observed in our experiment. In Figure 1, the “Global Recluster” curve exhibits more dramatic changes between rounds than other curves. Following the global re-clustering process, we need to determine the model each cluster should continue with. In our implementation, we align with previous works and use the average of the clients’ models within each cluster (Long et al., 2023; Zhang et al., 2024). When most clients remain in the same cluster, the new cluster models would be similar to the previous ones. But when there are significant changes in the clusters, either due to large drifts or bad centroid initialization, the new models will differ substantially and require additional time to stabilize. This leads to sudden accuracy drops as the new clusters warm up. In Figure 2(a), the x-axis denotes training rounds and the y-axis shows the mean accuracy we get with our selective global re-clustering (details in Section 3) subtracted by that of triggering global re-clustering by default. We observe that default global re-clustering leads to fluctuating test accuracy, which oscillates between meeting our approach’s accuracy and being as much as 5% worse.
Summary. Per-client adjustment fails in highly dynamic environments where many clients experience drifts. In contrast, global re-clustering may harm model accuracy because it increases the instability of cluster memberships and increases the model convergence times, especially during minor drifts. When cluster memberships remain largely stable, global re-clustering is not only excessive but also requires additional steps to warm up the new clusters (e.g., finding optimal cluster model parameters). On the other hand, when many clients experience drifts, moving individual clients can cause substantial shifts in cluster centers and miss the opportunity to create or remove clusters as distribution patterns evolve.
Based on the prior analysis, we opt for a design that starts with client individual migration and triggers global re-clustering selectively (Section 4 explains the detailed design). Our method is visualized as the red curve (“Move Ind.+Global Recluster”) in Figure 1. We retain a low heterogeneity value throughout training and achieve high model accuracy faster (Section  6).
3.2. Re-clustering All Drifted Clients
When we rerun the clustering algorithm every round, we can only easily recluster selected clients (Liu et al., 2023). Since selected clients already get the gradients or model parameters during training, it is easy to compare with the existing clusters and decide whether they experience drifts and need change. In contrast, since unselected clients do not report gradients, we cannot decide whether and where to move them.
However, as mentioned in Section 2.1, pedagogical FL settings typically select only a subset of clients to contribute in each round. When unselected clients experience data drifts and are still stuck in their previous cluster, they might be served by the cluster model trained on data of a different distribution. We plot the difference in clients’ mean test accuracy of re-clustering only selected clients versus all clients that experience data drifts every round in Figure 2(b). Re-clustering all drifted clients offers a 1.5%-4% higher accuracy in most rounds, which is a significant difference given that the final accuracy we achieve on FMoW is 52.4%. This trend demonstrates that only moving selected clients leads to imperfect clusters containing diverging but unselected clients.
Table 1. Heterogeneity of all clients and intra-cluster heterogeneity after gradient-based clustering (lower is better). At the initial stages, gradient-based clustering doesn’t sufficiently decrease heterogeneity due to the model being not well-trained. In contrast, label-based clustering provides a consistent decrease in heterogeneity.
Total Rounds Un-Clustered Gradient-Based Clustering Label-Based Clustering
100 1.81 1.82(+0.6%) 1.65(-8.8%)
200 1.80 1.82(+1.1%) 1.64(-8.9%)
500 1.78 1.71(-3.9%) 1.62(-9.0%)
1000 1.74 1.66(-4.6%) 1.56(-10.3%)
1500 1.76 1.63(-7.4%) 1.60(-9.1%)
Global re-clustering based on training outputs incur high overheads. Considering the suboptimal accuracy of only re-clustering the selected clients, our goal is to move all drifted clients promptly. One effective method to learn information from all clients for re-clustering is to train a global model with selected clients and then distribute it to all clients. For example, in FlexCFL, each client performs local training with the global model weights on their own data and reports gradients to the central server, where clustering happens (Duan et al., 2021). The coordinator can then re-cluster all clients according to the gradients.
This approach incurs high communication and computing overheads. For example, in our experiments training ResNet-18 (He et al., 2016) for image classification tasks, devices on average take 116.7 seconds to download the 11 million parameters and upload gradient coordinates. They also need an additional 50.4 seconds to run forward and backward propagation.
3.3. Label-based Instead of Gradient-based Clustering
Another issue with gradient-based re-clustering is that the clustering effectiveness is sensitive to the model quality. Table 1 presents the resulting mean client distance when we perform gradient-based clustering after training a global model for various numbers of rounds. The numbers in parentheses indicate the change in mean client distance relative to that of the global set. A negative value in the parentheses indicates that the generated clusters are overall less heterogeneous than the global set. When we perform gradient-based clustering with the global model at round 100 and round 200, the resulting clusters are more heterogeneous. As we postpone clustering to later rounds, gradient-based clustering achieves an increasingly larger reduction in mean client distance. This trend indicates that gradient-based clustering doesn’t perform well in earlier rounds, when the global model used to collect gradients from all clients hasn’t converged.
We propose using label distribution vectors for clustering (we refer to this approach as “label-based”). Label distribution vectors are available on all clients and enable us to promptly detect any drift by checking for distribution changes. Compared to gradient-based clustering, label-based doesn’t demand parameter or gradient transmission, doesn’t introduce additional computation tasks, and is not sensitive to the timing of client clustering. In Table 1, label-based clustering provides consistent heterogeneity reduction across rounds. Furthermore, as a label distribution vector typically has tens or hundreds of coordinates compared to millions in a full gradient, label-based clustering incurs significantly lower computational overhead. In our experiment of training ResNet-18 on the FMoW dataset, clustering all 302 clients using our label-based solution takes only 0.05 seconds while the gradient-based solution FlexCFL takes 37.92 seconds (note that FlexCFL already accelerates this process through dimensionality reduction using truncated Singular Value Decomposition).
4. Fielding Overview
Figure 3. Clustering overview of Fielding. Clients send distribution vectors to the coordinator; the coordinator moves drifted clients to the closest cluster; global re-clustering is triggered if any cluster center shifts by a distance larger than $\tau$.
To improve clustered FL performance, we develop Fielding, a clustered FL framework that efficiently reassigns clients after data drifts. As explained in Section 3, Fielding has three key design decisions: (1) reassign all drifted clients instead of only selected clients, (2) combine per-client adjustment and global re-clustering, and (3) cluster clients based on label distribution vectors.
4.1. System Overview
As discussed in Section 3.1, neither per-client adjustment nor global re-clustering works well across all drift scenarios. Therefore, we propose a combined approach that leverages both methods. Fielding employs a threshold $\tau$ to determine whether global re-clustering is necessary. When clients report data drifts, we first try re-clustering drifted clients individually by moving them to the cluster whose center is the closest to their new distribution vector. To ensure determinism across different orders of client movements, we do not update the cluster center throughout this per-client adjustment process. After individual movements, we recalculate the cluster centers and trigger global re-clustering over all clients only when any center moves by a distance larger than $\tau$. In our implementation, we set $\tau$ as one-third of the average distance between a pair of cluster centers, which is a value that works empirically well across evaluation settings.
Figure 3 illustrates the clustering mechanism of Fielding. When clients register with the centralized coordinator, they send their local data distribution vector, either during initial registration or following a recent data drift. Upon receiving this vector, the coordinator updates the client’s metadata and assigns the client to the nearest cluster. Before selecting participants for the upcoming training round, the coordinator finds the distance each cluster center has shifted. It then initiates a global clustering process if any cluster exceeds the threshold $\tau$. Otherwise, it finalizes the updated cluster membership.
Our clustering approach incurs additional communication costs, as clients are expected to upload their data distribution vector. However, this vector is small (4 bytes×$\times N$ where $N$ is the number of labels), and clients only need to send it when they first join or experience data drift. In Section 4.3, we prove the model convergence guarantees with label-based clustering.
4.2. Algorithmic Framework
1 Partition clients into clusters using Algorithm 3
2 Initialize cluster models $\mathbf{c}^{(1)},\ldots,\mathbf{c}^{(K)}$
3 for every iteration $t=0,\ldots,T-1$ do
4       Handle data drifts using Algorithm 2
5       Let $C_{1},\ldots,C_{K}$ be the clusters
6       Let $M$ be the number of machines sampled per iteration
7       for every cluster $k=1,\ldots,K$ in parallel do
8             for $R$ rounds do
9                   Sample a set $S$ of $M/K$ clients from $C_{k}$ based on the selection strategy
10                   for each client $i\in S$ in parallel do
11                         Initialize $\mathbf{x}_{i}=\mathbf{c}^{(k)}$
12                         for $L$ local iterations do
13                               $\mathbf{x}_{i}\leftarrow\mathbf{x}_{i}-\eta G^{(i)}_{t}(\mathbf{x}_{i})$
14
15                  $\mathbf{c}^{(k)}\leftarrow\operatorname*{avg}_{i\in S}\mathbf{x}_{i}$
16
17      report cluster models and client-to-cluster assignment
Algorithm 1 Clustered Federated Learning with data drift
1 for each drifting client $x$ do
2       Assign $x$ to the closest cluster center
3       Recompute the centers of the affected clusters
4Let $\theta$ be the average distance between the cluster centers
5 if some center moved by at least $\theta/3$ then
6       for each client $i$ do
7             Let $\mathbf{x}_{i}$ be the model corresponding to the $i$’th client’s cluster
8      Let $C_{1},\ldots,C_{K}$ be clusters returned by Algorithm 3
9       for each cluster $C_{k}$ do
10             $\mathbf{c}^{{(k)}}\leftarrow\operatorname*{avg}_{i\in C_{k}}\mathbf{x}_{i}$
11
Algorithm 2 Handling data drifts
1 Define the distance between clients as the $\ell_{1}$-distance between their label histograms
2 Choose $K$ with the largest silhouette score
Cluster the clients using the $K$-means clustering
Algorithm 3 Clustering
In this section, we describe our algorithm – presented in Algorithm 1 – in detail. Initially, Fielding clusters the clients using the $k$-means algorithm (Algorithm 3). Importantly, we measure distance between data distributions, which is computed as the $\ell_{1}$-distance between their histograms: that is, assuming we have $l$ labels, the distance between clients with label distribution $p_{1},\ldots,p_{l}$ and $q_{1},\ldots,q_{l}$ is defined as $\sum_{i=1}^{l}|p_{i}-q_{i}|$.
We assume that the algorithm is executed for $T$ global iterations, and during each iteration, the label distribution of each client might change, for instance by adding or removing data points. The first step of Fielding is to handle the data drifts (Algorithm 2). Our goal is, by the end of each iteration, to maintain a good model for each cluster, with respect to its clients’ data. In this step, each drifted client is assigned to the closest cluster. If this causes a significant change in cluster centers – namely, if there exists a cluster moving by more than $\tau/3$, where $\tau$ is the average distance between cluster centers – we use Algorithm 3 to recluster clients from scratch. This condition avoids frequent reclustering, which might require excessive resources and can adversely change the clients’ losses. Importantly, we recluster all the clients, not just available clients, which, as we show in Section 2, improves the accuracy.
For the global reclustering, we choose the number of clusters that gives the highest silhouette score. After reclustering, we compute new cluster models: for each client $i$, let $\mathbf{x}_{i}$ be its old cluster model, and then for each new cluster $C_{k}$, we define its model as an average of $\mathbf{x}_{i}$ for $i\in C_{k}$. That is, we average the old client models.
After the clusters are computed, we train the cluster models for $R$ rounds. During each round, we sample the clients from the cluster and run $L$ local iterations of gradient descent. After the local iterations, we set the cluster model as the average of the models of all participating clients.
4.3. Convergence
In Appendix A, we analyze the performance of our framework by bounding the average of clients’ loss functions at every iteration. Formally, for every cluster $C_{k}$, let $\mathbf{c}^{(k)}$ be the model corresponding to the cluster by the end of the iteration, and let $\mathbf{c}^{(k,*)}$ be the optimal model for this cluster, that is, the minimizer of $\sum_{i\in C_{k}}f^{(i)}_{t}$. Then, we bound the average loss of the clients with respect to the best possible cluster models:
(1) $\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in C_{k}}(f^{(i)}_{t}(\mathbf{c}^{(k)})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}))$
Our analysis has two ingredients, corresponding to two main stages in our algorithm: gradient descent steps and reclustering. For the gradient descent part, we use the standard convergence analysis of SGD, which for strongly convex function $f$ shows that for stochastic gradient descent update the following holds:
$\mathbb{E}\left[f(\mathbf{x}_{T})-f(\mathbf{x}^{*})\right]$ $\leq\left(1-\eta\mu\right)^{T}(f(\mathbf{x}_{0})-f(\mathbf{x}^{*}))+\frac{L\eta}{2\mu}\sigma^{2}$
The right-hand side has two terms, where the first term decays exponentially, and the second term corresponds to the total contribution of the stochastic noise. In our analysis, we similarly have the $(f(\mathbf{x}_{0})-f(\mathbf{x}^{*}))\left(1-\eta\mu\right)^{T}$ term which decays exponentially, and a non-vanishing term corresponding to the noise coming from the stochastic noise and the client sampling.
The major challenge in the analysis is bounding the effect of data drifts and reclustering. Data drifts change client datasets, hence changing their local objective functions. On the other hand, while reclustering is useful in the long run, its immediate effect on the objective (1) can potentially be negative (see Figure 2(b)) due to the mixing of models from different clusters.
To address these issues, we introduce natural assumptions. First, we assume that the distance between client representations – which in our algorithm correspond to label distributions – translates into the difference between their objective functions. Second, we assume that the effect of data drift is bounded: that is, the representation $r_{i}$ of client $i$ changes by at most $\delta$ for each data drift. Finally, we assume that the clients are clusterable; that is, $K$ clusters exist, so representations of clients within each cluster are similar.
Assumption A.
$f^{(i)}_{t}$ – local objective for client $i$ at iteration $t$ – is $\mu$-strongly convex and $L$-smooth, and we have access to a stochastic oracle with variance $\sigma^{2}$. There exists clustering such that representations inside each cluster are $\Delta$-close, the data drift changes representations by at most $\delta$, and the difference between the objective functions is at most the distance between representations multiplied by $\theta$.
Under these assumptions, the data drift can affect each client representation – and hence objective 1 – by at most a fixed value. Moreover, by the clustering assumption, the clients within each cluster have similar objective function; moreover, this holds both before and after the reclustering. This allows us to bound the increase in the loss function due to reclustering. In Appendix A, we prove the following theorem.
Theorem 4.1.
Let $N$ be the number of clients and $M$ be the total number of machines sampled per round. Let $\mathbf{x}^{*}$ be the minimizer of $f_{0}=\operatorname*{avg}_{i}f^{(i)}_{0}$. Let $\mathbf{c}^{(k,*)}_{t}$ be the minimizer for cluster $k$ at iteration $t$. Then, under A, for $\eta\leq 1/L$, for any iteration $t$ we have
$\frac{1}{N}\sum_{k\in[K]}\sum_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t+1})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$
$\leq(1-\eta\mu)^{TR}\left(f_{0}(\mathbf{x}_{0})-f_{0}(\mathbf{x}^{*})\right)$
$+\frac{L\eta}{2\mu}\left(\frac{\sigma^{2}+8L\theta\Delta}{M/K}+3\theta(\Delta+\delta)(1-\eta\mu)^{R}\right)$
That is, objective (1) has two terms: an exponentially decaying term corresponding to the initial loss, and a non-vanishing term corresponding to stochastic noise, as well as loss due to clustering and data drift.
5. Implementation
We implemented Fielding in Python on top of FedScale (Lai et al., 2022), a state-of-the-art open-source FL engine. We follow FedScale’s design of having one centralized coordinator and client-specific executors. Figure 4 illustrates the end-to-end workflow of Fielding. At the beginning of each training round, 1 client executors register with the coordinator to participate and report their local data distribution optionally. The coordinator’s cluster manager maintains clients’ distribution records and 2 moves clients to the closest cluster when their reported distribution drifts. After moving drifted clients individually, the cluster manager measures center shift distances and either triggers global re-clustering when any cluster center shifts significantly or finalizes clusters membership. In the case of global clustering, we use $K$-means clustering and determine $K$ using the silhouette method (Rousseeuw, 1987). To stabilize new clusters and avoid creating poor clusters due to bad centroids initialization (Ahmed et al., 2020), we reuse previous cluster centers when the target $K$ value is no greater than the number of existing clusters. The initial model of each newly created cluster is set as the average of its clients’ previous cluster model (see Algorithm 2).
Figure 4. The system architecture and workflow of Fielding. Blue arrows represent communication between clients and the coordinator. Black arrows indicate the workflow among different components on the same machine.
When client clustering is done, the coordinator notifies the client selector to 3 select a subset of clients to contribute to each cluster and 4 communicate the latest model parameters to the selected clients. 5 Upon receiving a set of parameters, client executors conduct the training process over local data and 6 upload the updated model parameters to the coordinator. Finally, 7 the model manager aggregates individual models and stores the new cluster models. This process happens iteratively until the clients’ mean test accuracy reaches a target value. Fielding also regularly creates checkpoints for the models, clients’ metadata, and cluster memberships for future fine-tuning and failure recovery. Fielding’s overheads are mainly determined by the number and size of distribution vectors we need to re-cluster. In our largest evaluation setting where we train with 5078 clients on a dataset with 100 labels, per-client adjustment takes 2.0 seconds and global re-clustering takes 15.6 seconds on average. Storing the latest reported distribution vector for each client consumes $5078\times 100\times 4$B $\approx 1.9$MB of memory.
6. Evaluation
We evaluate Fielding’s effectiveness by running four FL tasks. We also demonstrate Fielding’s compatibility with various client selection strategies and aggregation algorithms. Our end-to-end results highlight that
  • Fielding improves the final test accuracy by 1.9%-5.9% and is more stable than other clustered FL works when facing real-world data drifts.
  • Fielding works well with complementary FL optimizations and offers consistent accuracy improvements across different settings.
  • Fielding is robust to malicious clients who report corrupted distribution vectors and accelerates model convergence when clients are less heterogeneous.
Figure 5. Time to accuracy (TTA) comparison over four tasks.
Environment. We emulate large-scale FL training with four NVIDIA A100 GPUs. We use FedScale’s device datasets which provide realistic device computing and network capacity numbers while following the standardized training configurations in the original paper (Lai et al., 2022). Per-device computation time estimation is based on device computing speed, batch size, and number of local iterations. And we estimate the communication time as the volume of downloaded and uploaded data divided by the network bandwidth.
Datasets and Models. We use a satellite image dataset (FMoW (Christie et al., 2018)) and two video datasets (Waymo Open (Sun et al., 2020) and Cityscapes (Cordts et al., 2016)) that include natural data drifts. For FMoW, we keep images taken after January 1, 2015, and create one client for each UTM zone. Two training rounds correspond to one day in this dataset. For Waymo Open and Cityscapes, we use video segments pre-processed by Ekya (Bhardwaj et al., 2022). We create one client per camera per Waymo Open segment (212 clients in total), and one client per 100 consecutive frames per Cityscapes segment (217 clients in total). As these two datasets have video frame IDs within segments but lack global timestamps, we sort the samples on each client by frame ID and then partition the frames into 10 intervals. We stream in one data interval every 20 training rounds. We train ResNet-18 (He et al., 2016) on these three real-world datasets.
To demonstrate how Fielding behaves in highly dynamic settings, we construct an artificial trace using the FedScale Open Images (Kuznetsova et al., 2020) benchmark. We find the top 100 classes in terms of number of samples and keep clients whose local data labels include at least 10 of the top 100 classes (5078 clients in total). Each client randomly and independently partitions all local data labels into 10 buckets. Each bucket then contains all local samples of labels mapped to it. We stream in one bucket every 50 training rounds. We train ShuffleNet v2 (Ma et al., 2018) on Open Images. On all four datasets, clients retain samples they received over the last 100 rounds. The details of experiment settings and training parameters are included in Appendix B.
Baselines. Our non-clustering baseline trains one global model by randomly selecting a subset of participants from all available clients every round. To compare with other clustered FL works with data drift handling measures, we use Auxo (Liu et al., 2023) as the continuous re-clustering baseline and FlexCFL (Duan et al., 2021) as the individual movement baseline. We employ FedProx (Li et al., 2020b), a federated optimization algorithm that tackles client heterogeneity, for all approaches.
Metrics. Our main evaluation metrics are Time-to-Accuracy (TTA) and final test accuracy. We define TTA as the training time required to achieve the target accuracy, which is the average test accuracy when training a single global model for all clients. We report the final average test accuracy across diverse settings to demonstrate the sensitivity and robustness of Fielding.
6.1. End-to-end Training Performance
Figure 5 shows that Fielding improves clients’ average test accuracy by 5.9%, 1.9%, and 2.1% on FMoW, Cityscapes, and Waymo Open respectively. When considering the global model’s final test accuracy as the target accuracy, Fielding offers a 1.27$\times$, 1.16$\times$, and 2.61$\times$ training speed up respectively (we define the moment when Fielding’s accuracy starts invariably exceeding the target accuracy as the point at which it achieves the target accuracy). On the highly heterogeneous and dynamic Open Images trace, Fielding achieves a significant accuracy improvement of 26%.
Although Auxo exhibits comparable or better performance than the global model baseline, the benefits are limited: accuracy improvement up to 0.9% and convergence speedup up to 1.05$\times$. These results are less impressive than those reported in the original paper on static datasets, suggesting a decline in clustering effectiveness. This degradation stems from re-clustering only the selected participants after each round, which neglects clients that have drifted but remain unselected. Additionally, this approach can lead to sudden accuracy drops on the biased Waymo Open dataset, where over 80% samples are cars, as selecting diverging clients can cause dramatic model weight changes. FlexCFL fails to improve model accuracy or convergence speed on all three real-world datasets. This is because FlexCFL migrates clients individually without adjusting the number of clusters. As we observe in Figure 1, such an approach will eventually make clusters as heterogeneous as a global set containing all clients.
(a) Oort selection
(b) Distance-based selection
Figure 6. Fielding with different client selection strategies on FMoW dataset.
6.2. Compatibility with Other Optimizations
To demonstrate that Fielding is agnostic to client selection and model aggregation strategies, we evaluate Fielding’s performance on FMoW together with various complementary optimizations. In Figure 6 and Figure 7, we use the algorithm name alone to denote the baseline where we train one global model and use “+ Fielding ” to denote running the algorithm atop Fielding.
We use the Oort selection algorithm (Lai et al., 2021) and a distance-based algorithm prioritizing clients closer to the distribution center as client selection examples. As shown in Figure 6, Fielding improves the final average test accuracy by 6.0% and 4.5% while reaching the target accuracy 2.62×2.62$\times$2.62 × and 1.17×1.17$\times$1.17 × faster. Note that Oort selection actively incorporates clients otherwise filtered out due to long response time, leading to a longer average round time.
We also show that Fielding works well with aggregation algorithm FedYogi (Reddi et al., 2021) and q-FedAvg (Li et al., 2020c) in Figure 7. Fielding improves the final accuracy by 5.9% and 4.9% while giving a 1.34×1.34$\times$1.34 × and 1.25×1.25$\times$1.25 × speedup respectively.
6.3. Robustness and Sensitivity Analysis
Malicious clients. We study the impact of having malicious clients who intentionally report wrong label distribution vectors. We simulate such behavior by randomly selecting a subset of clients who permute the coordinates of their distribution vectors when registering with the coordinator. As reported in Figure 10, Fielding consistently outperforms the baseline global model accuracy across different percentages of malicious clients.
(a) FedYogi
(b) q-FedAvg
Figure 7. Fielding with different FL algorithms on FMoW dataset.
Figure 8. Fielding is robust to malicious clients.
Figure 9. Fielding offers gains across heterogeneity degrees.
Figure 10. Fielding offers gains across heterogeneity degrees.
Varying data heterogeneity. Our analysis on Fielding’s sensitivity to data heterogeneity degrees is inspired by the idea of improving training performance on non-IID data through a small shared dataset (Zhao et al., 2018). We inject new training samples by creating three datasets shared with all clients: one sample for each of the least represented 50% labels, one sample for each label, and two samples for each label. A larger shared set implies a smaller degree of heterogeneity. As shown in Figure  10, both Fielding and the baseline global model benefit from this sharing approach, with Fielding improving the final average test accuracy by 3.2% to 4.4%.
Static data. Finally, we demonstrate that Fielding still improves model convergence when clients have static local data and no drifts happen. We rerun the experiment on the FMoW dataset with all data samples available throughout the training. As shown in Figure 10, Fielding offers the largest gain and improves the final average test accuracy by 9.2%. Auxo and FlexCFL achieve an improvement of 3.9% and 5.6% respectively. Note that Auxo experiences an accuracy drop towards the end of training. This is likely due to Auxo re-clustering all selected clients after each round by default, causing unnecessary cluster membership changes when the data is completely static.
7. Related Work and Discussion
Clustered FL Previous clustering solutions (Ghosh et al., 2020; Duan et al., 2021; Liu et al., 2023; Zhang et al., 2024; Jothimurugesan et al., 2023) have achieved success at bounding intra-cluster heterogeneity and improving model performance. However, their improvements are compromised by real-world data drifts as they either assume stable environments or handle drifts insufficiently.
Heterogeneity-aware FL. Recent works have tackled heterogeneity challenges through client selection, workload scheduling, and update corrections. Oort considers both statistical and system utility and designs an exploration-exploitation strategy for client selection (Lai et al., 2021). PyramidFL adjusts the number of local iterations and parameter dropouts to optimize participants’ data and system efficiency (Li et al., 2022b). DSS-Edge-FL dynamically determines the size of local data used for training and selects representative samples, optimizing resource utilization while considering data heterogeneity (Serhani et al., 2023). MOON corrects local updates on clients by maximizing the similarity between representations learned by local models and the global model (Li et al., 2021).
Drifts-aware FL. Adaptive-FedAvg is an FL algorithm that adapts client learning rate based on the model variability between consecutive rounds (Canonaco et al., 2021). CDA-FedAvg detects concept drifts and extends FedAvg for learning under such drifts. It defines a short-term and a long-term memory for each client and applies rehearsal using data in the lone-term memory when drifts happen (Casado et al., 2022). Master-FL proposes a multi-scale algorithmic framework that instructs clients to train over multiple time horizons with a suitable learning rate for each time chunk (Ganguly and Aggarwal, 2023).
Fielding focuses on dynamic client clustering that is robust against potential data drifts. It is agnostic to other FL optimizations and should work seamlessly with the FL aggregation, selection, and scheduling algorithms above.
8. Conclusion
We have presented Fielding, a novel clustering-based Federated Learning framework designed to handle data drifts with low overhead. Fielding addresses the challenges of data drift by re-clustering clients based on their label distribution vectors as their data distributions change, dynamically adjusting the number of clusters, and stabilizing responses to minor data shifts. Fielding is robust against malicious clients and varying degrees of heterogeneity, and it is compatible with various client selection and aggregation algorithms. In Fielding, clients register with a centralized coordinator that updates their metadata and assigns them to the nearest cluster based on their local data distribution. Before each training round, the coordinator determines how much each cluster center has shifted. If any cluster exceeds a predefined threshold, a global re-clustering process is initiated; otherwise, the updated cluster memberships are finalized. Experimental results demonstrate that Fielding improves the final test accuracy by 1.9% to 5.9% compared to existing methods and provides greater stability when facing real-world data drifts across different settings.
There are potential directions to further enhance Fielding. Our current approach requires labeled data, which may not be available in all contexts. Additionally, clients actively report their label distribution vectors to the coordinator, which introduces latent privacy concerns. In future work, we will focus on alternatives, such as inferring distribution vectors from gradients.
Appendix A Proof of Convergence
1 input: the number of clients $K$, the number of data drift events $T$, the number of sampled clients per round $M$, the number of rounds per data drift event $R$
2 Initialize cluster models $\mathbf{c}^{(1)}_{0},\ldots,\mathbf{c}^{(K)}_{0}$ with the same model $\mathbf{x}_{0}$
3 for $t=0,\ldots,T-1$ do
4       for each client $i$ do
5             Adopt data drift for client $i$
6             Reassign client $i$ to the closest cluster
7             Let $\mathbf{x}_{i}$ be the model corresponding to the cluster containing client $i$
8      Let $C_{1},\ldots,C_{K}$ be the $K$-center clustering of all clients based on representations from model $\tilde{\mathbf{x}}_{\tau}$
9       for $k=1,\ldots,K$ do
10             $\tilde{\mathbf{c}}^{(k)}_{0}\leftarrow\operatorname*{avg}_{i\in C_{k}}\mathbf{x}_{i}$
11      for $\tau=0,\ldots,R-1$ do
12             for $k=1,\ldots,K$ do
13                   Sample subset $S$ from $C_{k}$ of size $M/K$ independently with replacement
14                   $\tilde{\mathbf{c}}^{(k)}_{\tau+1}\leftarrow\tilde{\mathbf{c}}^{(k)}_{\tau}-\eta\operatorname*{avg}_{i\in S}G^{(i)}_{t}(\tilde{\mathbf{c}}^{(k)}_{\tau})$
15
16      for $k=1,\ldots,K$ do
17             $\mathbf{c}^{(k)}_{t+1}\leftarrow\tilde{\mathbf{c}}^{(k)}_{R}$
18
Algorithm 4 Clustered Federated Learning with Data Drift
In this section, we analyze the performance of our framework in the case when the number of clusters $K$ is known and the number of local iterations is $1$. The analysis can be modified to handle local iterations and other variations based on the previous work, see e.g. Li et al. (2020a). In the algorithm, we recluster clients after every data drift event and use the $k$-center clustering instead of $k$-means clustering to simplify the proof.
We present the algorithm in Algorithm 4. After every data drift event, we first accept data changes for each client. After that, we recluster the clients, and assign to each cluster a model by averaging models of all clients in the cluster. After that, for several rounds, we perform standard Federated Averaging updates on each cluster: we sample $M$ clients, compute the stochastic gradient for each client, and update the cluster model with the sampled gradients.
Assumption B (Objective functions).
Let $f^{(i)}_{t}$ be the local function at each client $i$ at data drift event $t$. Then $f^{(i)}_{t}$ is $\mu$-strongly convex and $L$-smooth (has an $L$-Lipschitz gradient) for all $i$ and $t$.
The following are the standard assumptions on stochastic gradient.
Assumption C (Stochastic gradients).
For each client $i$, at every data drift event $t$, we have access to stochastic gradient oracle $G^{(i)}_{t}$:
  • $\mathbb{E}\left[G^{(i)}_{t}(\mathbf{x})\right]=f^{(i)}_{t}(\mathbf{x})$ for all $\mathbf{x}$;
  • $\mathbb{E}\left\|G^{(i)}_{t}(\mathbf{x})-f^{(i)}_{t}(\mathbf{x})\right\|^{2}\leq\sigma^{2}$ for all $\mathbf{x}$.
The next assumption connects the objective value with client representations. Client representations can take various forms, and in our paper, we assume that the representation of the cluster is its label distribution.
Assumption D (Representation).
For a metric space $(X,\rho)$, let $r^{(i)}_{t}\in X$ be a representation of client $i$ at time $t$. Then there exists a constant $\theta$ such that $\left|f^{(i)}_{t}(\mathbf{x})-f^{(j)}_{t}(\mathbf{x})\right|\leq\theta\cdot\rho\left(r^{(i)}_{t},r^{(j)}_{t}\right)$ for all $i,j,t$.
Clearly, if data on clients can change arbitrarily at every data drift event, in general, there is no benefit in reusing previous iterates. Hence, our next assumption bounds how much the client changes after the drift.
Assumption E (Data drift).
At every data drift event, the representation of each client changes by at most $\delta$: $\rho\left(r^{(i)}_{t},r^{(i)}_{t+1}\right)\leq\delta$ for all $i$, $t$.
Finally, we assume that the clients are clusterable: there exist $K$ clusters so that within each cluster all points have similar representation.
Assumption F (Clustering).
At every data drift event, the clients can be partitioned into $K$ clusters $C_{1},\ldots,C_{K}$ so that, for any $k\in[K]$ and any $i,j\in C_{k}$, we have $\rho\left(r^{(i)}_{t},r^{(j)}_{t}\right)\leq\Delta$.
For completeness, we present the SGD convergence analysis for strongly convex functions. The following analysis is well-known, but, compared with the most standard analysis, it works directly with the loss functions instead of distances to the optimum, thus avoiding losing the condition number in the following bounds.
Lemma A.1.
Let $f$ be $L$-smooth and $\mu$-strongly convex. Let $g$ be an unbiased stochastic gradient oracle of $f$ with variance $\sigma^{2}$. Let $\mathbf{x}^{*}$ be the minimizer of $f$. Then, for the stochastic gradient update rule $\mathbf{x}_{t+1}\leftarrow\mathbf{x}_{t}-\eta g(\mathbf{x}_{t})$ with $\eta\leq 1/L$, we have
$\mathbb{E}\left[f(\mathbf{x}_{T})-f(\mathbf{x}^{*})\right]$ $\leq\left(1-\eta\mu\right)^{T}(f(\mathbf{x}_{0})-f(\mathbf{x}^{*}))+\frac{L\eta}{2\mu}\sigma^{2}.$
Proof.
Let $\mathbb{E}_{t}\left[\cdot\right]$ be expectation conditioned on $\mathbf{x}_{t}$. By the Descent Lemma, we have:
$\mathbb{E}_{t}\left[f(\mathbf{x}_{t+1})\right]$ $\leq f(\mathbf{x}_{t})+\mathbb{E}_{t}\left\langle\nabla f(\mathbf{x}_{t}),\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\rangle+\frac{L}{2}\mathbb{E}_{t}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}$
$=f(\mathbf{x}_{t})+\mathbb{E}_{t}\left\langle\nabla f(\mathbf{x}_{t}),-\eta g(\mathbf{x}_{t})\right\rangle+\frac{L}{2}\mathbb{E}_{t}\left\|\eta g(\mathbf{x}_{t})\right\|^{2}$
$=f(\mathbf{x}_{t})-\eta\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}+\frac{L\eta^{2}}{2}\left(\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}+\sigma^{2}\right)$
$\leq f(\mathbf{x}_{t})-\frac{\eta}{2}\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}+\frac{L\eta^{2}}{2}\sigma^{2},$
where in the last inequality we used $\eta\leq\frac{1}{L}$. Using the fact that $f$ is $\mu$-strongly-convex, we have $\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}\geq 2\mu(f(\mathbf{x}_{t})-f(\mathbf{x}^{*}))$, giving
$\mathbb{E}_{t}\left[f(\mathbf{x}_{t+1})\right]$ $\leq f(\mathbf{x}_{t})-\eta\mu(f(\mathbf{x}_{t})-f(\mathbf{x}^{*}))+\frac{L\eta^{2}}{2}\sigma^{2}.$
Subtracting $f(\mathbf{x}^{*})$ from both parts, we have
$\mathbb{E}_{t}\left[f(\mathbf{x}_{t+1})-f(\mathbf{x}^{*})\right]$ $\leq\left(1-\eta\mu\right)(f(\mathbf{x}_{t})-f(\mathbf{x}^{*}))+\frac{L\eta^{2}}{2}\sigma^{2}.$
By telescoping, taking the full expectation, and using $\sum_{i=0}^{\infty}\left(1-\eta\mu\right)^{i}=\frac{1}{\eta\mu}$, we get
$\mathbb{E}\left[f(\mathbf{x}_{T})-f(\mathbf{x}^{*})\right]$ $\leq\left(1-\eta\mu\right)^{T}(f(\mathbf{x}_{0})-f(\mathbf{x}^{*}))+\frac{L\eta}{2\mu}\sigma^{2}.\qed$
The above result shows how the objective improves with each round: namely, after $R$, the non-stochastic part of the loss improves by a factor which depends on $R$ exponentially. However, each data drift and each reclustering might potentially hurt the objective. Our next result bounds their effect on the loss.
Lemma A.2.
Let $N$ be the number of clients, and $t+1$ be a fixed data drift event. Let $C_{1},\ldots,C_{k}$ be clusters $\mathbf{c}_{1},\ldots,\mathbf{c}_{K}$ be cluster models, and $\mathbf{c}^{*}_{1},\ldots,\mathbf{c}^{*}_{K}$ be the optimal cluster models at the end of processing data drift event $t$. Similarly, let $\bar{C}_{1},\ldots,\bar{C}_{\ell}$ be clusters $\bar{\mathbf{c}}_{1},\ldots,\bar{\mathbf{c}}_{K}$ be cluster models, and $\bar{\mathbf{c}}^{*}_{1},\ldots,\bar{\mathbf{c}}^{*}_{K}$ be the optimal cluster models immediately after reclustering. Then, under Assumptions B-D, the following holds:
$\frac{1}{N}\sum_{\ell\in[K]}\sum_{i\in\bar{C}_{\ell}}\left(f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell})-f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell}^{*})\right)\leq\frac{1}{N}\sum_{k\in[K]}\sum_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}_{k})-f^{(i)}_{t}(\mathbf{c}_{k}^{*})\right)+3\theta(\Delta+\delta)$
Proof.
For a client $i$, let $k$ and $\ell$ be such that $i\in C_{k}\cap\bar{C}_{\ell}$. We rewrite $f^{(i)}_{t}(\bar{\mathbf{c}}_{\ell})-f^{(i)}_{t}(\bar{\mathbf{c}}_{\ell}^{*})$ as
$f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell})-f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell}^{*})=(f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell})-f^{(i)}_{t+1}(\mathbf{c}_{k}^{*}))+(f^{(i)}_{t+1}(\mathbf{c}_{k}^{*})-f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell}^{*}))$
and bound each term separately.
Bounding the first term
Let $\mathbf{x}_{j}$ be the model of client $j$ before reclustering, i.e. $\mathbf{x}_{j}=\mathbf{c}_{k^{\prime}}$ where $j\in C_{k^{\prime}}$. Then,
(2) $f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell})=f^{(i)}_{t+1}\Bigl{(}\operatorname*{avg}_{j\in\bar{C}_{\ell}}\mathbf{x}_{j}\Bigr{)}\leq\operatorname*{avg}_{j\in\bar{C}_{\ell}}f^{(i)}_{t+1}(\mathbf{x}_{j}),$
where the last inequality follows by convexity of $f^{(i)}_{t}$. By F, since all clients $\bar{C}_{\ell}$ belong to the same cluster, their representations differ by at most $\Delta$, and hence by Assumption D their local objectives differ by at most $\theta\Delta$. Therefore,
$\operatorname*{avg}_{j\in\bar{C}_{\ell}}f^{(i)}_{t+1}(\mathbf{x}_{j})\leq\operatorname*{avg}_{j\in\bar{C}_{\ell}}(f^{(j)}_{t+1}(\mathbf{x}_{j})+\theta\Delta)=\operatorname*{avg}_{j\in\bar{C}_{\ell}}f^{(j)}_{t+1}(\mathbf{x}_{j})+\theta\Delta$
Summing over all $i\in\bar{C}_{\ell}$, we get
$\sum_{i\in\bar{C}_{\ell}}(\operatorname*{avg}_{j\in\bar{C}_{\ell}}f^{(i)}_{t+1}(\mathbf{x}_{j})+\theta\Delta)=\sum_{j\in\bar{C}_{\ell}}(f^{(j)}_{t+1}(\mathbf{x}_{j})+\theta\Delta)$
By definition, for each $i\in C_{k}$ we have $\mathbf{x}_{i}=\mathbf{c}_{k}$. So,
$\sum_{\ell\in[K]}\sum_{i\in\bar{C}_{\ell}}f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell})\leq\sum_{k\in[K]}\sum_{i\in C_{k}}f^{(i)}_{t+1}(\mathbf{c}_{k})$
Hence, the sum of the first terms in Equation 2 over all $i$ can be bounded as
$\sum_{k\in[K]}\sum_{i\in C_{k}}(f^{(i)}_{t+1}(\mathbf{c}_{k})-f^{(i)}_{t+1}(\mathbf{c}_{k}^{*})+\theta\Delta)\leq\sum_{k\in[K]}\sum_{i\in C_{k}}(f^{(i)}_{t}(\mathbf{c}_{k})-f^{(i)}_{t}(\mathbf{c}_{k}^{*})+\theta(\Delta+2\delta)),$
where we used Assumptions D and E to bound the change of each objective after the data drift as $\theta\delta$.
Bounding the second term
Let $\mathbf{x}_{t}^{(i,*)}$ be the minimizer of $f^{(i)}_{t}$ and $\mathbf{x}_{t+1}^{(i,*)}$ be the minimizer of $f^{(i)}_{t+1}$. Then, by Assumptions D and F we have
$f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell}^{*})$ $\geq f^{(i)}_{t+1}(\mathbf{x}_{t+1}^{(i,*)})$ ($\mathbf{x}_{t+1}^{(i,*)}$ is the minimizer of $f^{(i)}_{t+1}$)
$\geq f^{(i)}_{t}(\mathbf{x}_{t+1}^{(i,*)})-\theta\delta$ (Assumptions D and E)
$\geq f^{(i)}_{t}(\mathbf{x}_{t}^{(i,*)})-\theta\delta$ ($\mathbf{x}_{t}^{(i,*)}$ is the minimizer of $f^{(i)}_{t}$)
$\geq f^{(i)}_{t}(\mathbf{c}_{k}^{*})-\theta(\Delta+\delta),$
where the last inequality holds since otherwise $f^{(i)}_{t}(\mathbf{c}_{k}^{*})>f^{(i)}_{t}(\mathbf{x}_{t}^{(i,*)})+\theta\Delta$, which is impossible since $\mathbf{c}_{k}^{*}$ is the minimizer of $\operatorname*{avg}_{j\in C_{k}}f^{(j)}_{t}$ and $\operatorname*{avg}_{j\in C_{k}}f^{(j)}_{t}(\mathbf{x}_{t}^{(i,*)})\leq f^{(i)}_{t}(\mathbf{x}_{t}^{(i,*)})+\theta\Delta$ by Assumptions D and 3.
Finally, we have $|f^{(i)}_{t}(\mathbf{c}_{k}^{*})-f^{(i)}_{t+1}(\mathbf{c}_{k}^{*})|<\theta\delta$ by Assumptions D and E. Combining the bounds concludes the proof. ∎
Combining these results leads to our main convergence result.
Theorem A.3.
Let $N$ be the number of clients, and let their objective functions satisfy Assumptions B-D. Let $M$ be the total number of machines sampled per round. Let $\mathbf{x}^{*}$ be the minimizer of $f_{0}=\operatorname*{avg}_{i}f^{(i)}_{0}$. Let $\mathbf{c}^{(k,*)}_{t}$ be the minimizer for cluster $k$ at data drift event $t$. Then, for $\eta\leq 1/L$, for any data drift event $t$ of Algorithm 4 we have
$\frac{1}{N}\sum_{k\in[K]}\sum_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t+1})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)\leq(1-\eta\mu)^{TR}\left(f(\mathbf{x}_{0})-f(\mathbf{x}^{*})\right)+\frac{L\eta}{2\mu}\left(\frac{\sigma^{2}+8L\theta\Delta}{M/K}+3\theta(\Delta+\delta)(1-\eta\mu)^{R}\right)$
Proof.
First, we express the objective in terms of objectives for individual clusters:
$\frac{1}{N}\sum_{k\in[K]}\sum_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)=\frac{1}{N}\sum_{k\in[K]}|C_{k}|\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$
We then analyze the convergence in each cluster. For any $L$-smooth function $h$ with minimizer $\mathbf{x}^{*}$, it holds that $\|\nabla h(\mathbf{x})\|\leq\sqrt{2L(h(\mathbf{x})-h(\mathbf{x}^{*}))}$. First, note that, inside each cluster, the objective functions differ by at most $\theta\Delta$. Hence, for any two clients $i$ and $j$ from the same cluster, defining $h=f^{(i)}_{t}-f^{(j)}_{t}$, by $2L$-smoothness of $h$, for any $\mathbf{x}$ we have
$\|\nabla h(\mathbf{x})\|\leq\sqrt{4L(h(\mathbf{x})-h(\mathbf{x}^{*}))}\leq\sqrt{8L\theta\Delta}$
Hence, sampling clients uniformly from the cluster introduces variance at most $8L\theta\Delta$. Since we sample $M/K$ clients from a cluster, the total variance – including the stochastic variance – is at most $\frac{\sigma^{2}+8L\theta\Delta}{M/K}$.
Next, by Lemma A.2, the total objective increases by at most $3\theta(\Delta+\delta)$ after reclustering. Since during $R$ rounds the non-stochastic part of each objective decreases by a factor of $(1-\eta\mu)^{R}$, we have
$\frac{1}{N}\sum_{k\in[K]}|C_{k}|\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t+1}(\mathbf{c}^{(k)}_{t+1})-f^{(i)}_{t+1}(\mathbf{c}^{(k,*)}_{t+1})\right)$
$\leq\frac{1}{N}\sum_{k\in[K]}|C_{k}|\left(\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)+3\theta(\Delta+\delta)\right)(1-\eta\mu)^{R}+\sum_{i=0}^{R}\left(1-\eta\mu\right)^{i}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$
$=(1-\eta\mu)^{R}3\theta(\Delta+\delta)+(1-\eta\mu)^{R}\frac{1}{N}\sum_{k\in[K]}|C_{k}|\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)+\sum_{i=0}^{R}\left(1-\eta\mu\right)^{i}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$
By Lemma A.1, the last (stochastic) term accumulates over all data drift events as $\frac{L\eta}{2\mu}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$. By a similar reasoning, $3\theta(\Delta+\delta)$ accumulates as $\frac{L\eta}{2\mu}\cdot 3\theta(\Delta+\delta)(1-\eta\mu)^{R}$. And finally, $f(\mathbf{x}_{0})-f(\mathbf{x}^{*})$ term is multiplied by $(1-\eta\mu)^{R}$ at every data drift event, giving factor of $(1-\eta\mu)^{TR}$ over $T$ data drift events. ∎
Appendix B Experiment settings
We conduct experiments on public datasets Functional Map of the World (FMoW) (Christie et al., 2018), Cityscapes (Cordts et al., 2016), Waymo Open (Sun et al., 2020), and Open Images (Kuznetsova et al., 2020). In Table 2, we list the total number of clients we create on each dataset and the number of rounds in between clients getting new samples (e.g., as mentioned in Section 6, on Cityscapes we partition each client’s data into 10 intervals and stream in one interval every 30 rounds; so the rounds between new data arrivals are 30). In Table 3, we specify the training parameters including the total number of training rounds, learning rate, batch size, number of local steps, and the total number of clients selected to contribute in each round.
Table 2. Dataset configurations.
Dataset Number of Clients Rounds Between New Data Arrivals
FMoW 302 2
CityScapes 217 30
Waymo Open 212 20
Open Images 5078 50
Table 3. Training parameters.
Dataset Total Rounds Learning Rate Batch Size Number of Local Steps Clients Selected per Round
FMoW 2000 0.05 20 20 50
CityScapes 200 0.001 20 20 20
Waymo Open 100 0.001 20 20 20
Open Images 400 0.05 20 20 200

### v2

@@
-Federated Learning Clients Clustering with Adaptation to Data Drifts
-Abstract.
-Federated Learning (FL) enables deep learning model training across edge devices and protects user privacy by retaining raw data locally.
-Data heterogeneity in client distributions slows model convergence and leads to plateauing with reduced precision.
-Clustered FL solutions address this by grouping clients with statistically similar data and training models for each cluster.
-However, maintaining consistent client similarity within each group becomes challenging when data drifts occur, significantly impacting model accuracy.
-In this paper, we introduce Fielding, a clustered FL framework that handles data drifts promptly with low overheads.
-Fielding detects drifts on all clients and performs selective label distribution-based re-clustering to balance cluster optimality and model performance, remaining robust to malicious clients and varied heterogeneity degrees.
-Our evaluations show that Fielding improves model final accuracy by 1.9%-5.9% and reaches target accuracies 1.16$\times$-2.61$\times$ faster.
-copyright: nonedoi: isbn: price:
-1. Introduction
-Federated Learning (FL) is a distributed machine learning paradigm that enables multiple clients to train one model collaboratively while retaining raw data locally (Bonawitz et al., 2019; McMahan et al., 2017).
-As FL circumvents the communication costs of data movement and mitigates privacy leakage (Lyu et al., 2024; Huba et al., 2022), it has drawn attention across various fields and seen adoption in applications such as Google Keyboard (Yang et al., 2018), energy consumption forecasting (Abdulla et al., 2024), and medical image processing (Zhou et al., 2023).
-The main challenge in federated learning is data heterogeneity.
-Since federated learning involves many devices of diverse settings, there exists substantial dataset size and distribution divergences between clients from distant geographical locations and with distinct personal habits (Li et al., 2022a; Kim et al., 2024).
-Such data heterogeneity leads to slow model convergence and plateauing with degraded precision (Li et al., 2020a; Zhao et al., 2018; Sattler et al., 2020).
-A wide range of clustering mechanisms have been proposed to group clients with similar data distribution into more homogeneous clusters and train one model for each cluster (Ghosh et al., 2020; Duan et al., 2021; Liu et al., 2023; Jothimurugesan et al., 2023).
-Data drift occurs frequently in federated learning systems.
-It is defined as temporal changes in the label distribution (i.e., the proportion of samples assigned to each label) of clients’ local datasets.
-Such drifts arise from varying sampling rates, hardware configurations, sensor modalities, and limitations in client storage capacities (Chen et al., 2020; Nandi and Xhafa, 2022; Marfoq et al., 2023).
-Existing clustering mechanisms, however, are susceptible to data drift.
-Prior works address this issue but often either inadequately adjust clusters or incur prohibitive costs.
-IFCA and FlexCFL (Ghosh et al., 2020; Duan et al., 2021) adjust clusters by relocating shifted clients while maintaining a fixed number of clusters, operating under the assumption of low drift magnitude.
-Auxo (Liu et al., 2023) re-clusters clients participating in the current training round using gradient information; however, since only a subset of clients is selected each round, it delays adjustments for drifted clients until they are chosen for training.
-FedDrift (Jothimurugesan et al., 2023) re-clusters all clients and adaptively determines the number of clusters by transmitting all cluster models to every client each round to obtain training loss on local data, leading to significant communication and computation costs.
-In this work, we propose Fielding, a novel clustering-based FL framework that retains cluster client similarity despite frequent data drifts.
-Fielding focuses on re-clustering clients as data drifts, dynamically adjusting the number of clusters while stabilizing responses to minor data drifts.
-Fielding combines per-client migration with selective global clustering to capture significant distribution patterns without compromising accuracy.
-We assume the existence of a coordinator, typically deployed on a powerful central server to manage synchronization and advance global model training (Bonawitz et al., 2019; Bagdasaryan et al., 2020).
-Our solution handles all drifted clients by clustering them according to their local distribution vector, a piece of information readily available all the time.
-We present a theoretical framework supporting this approach.
-Fielding is designed to be lightweight; it introduces no additional computation on the client side and requires clients only to transmit their distribution vector (an array with up to a few thousand values) when they register with the coordinator.
-We implement Fielding by extending the widely adopted FL engine FedScale (Lai et al., 2022) to support streaming data and our clustering mechanisms.
-We evaluated Fielding with up to 5078 clients on four image streaming traces with data drifts.
-Compared to existing approaches, Fielding improves model accuracy by 1.9% to 5.9% and achieves the target accuracy 1.16$\times$ to 2.61$\times$ faster.
-We also show that Fielding is compatible with various client selection and aggregation algorithms, and robust to malicious clients and varied heterogeneity degrees.
-Our main contributions are:
-  • We introduce Fielding, a clustered FL framework that handles data drifts promptly with low overheads.
-  • We show that Fielding imposes minimal overhead by eliminating additional client-side computations and reducing communication costs through the use of readily available data, and we provide a theoretical framework to support our approach.
-  • We extend the FedScale FL engine to implement Fielding and demonstrate its efficacy on four real-world datasets with up to 5,078 clients, achieving up to 5.9% accuracy improvement and faster convergence compared to existing methods.
-  • We show that Fielding is compatible with various client selection and aggregation algorithms and is robust against malicious clients and different degrees of data heterogeneity.
-2. Background
-In this section, we provide the background of clustered federated learning and the impacts of data drifts.
-2.1. Clustered Federated Learning
-FL deployment often involves hundreds to thousands of available clients participating in the training (Bonawitz et al., 2019).
-For example, Google sets 100 as the target number of clients for a round of training when improving keyboard search suggestions (Yang et al., 2018) and (Huang et al., 2019) have developed mortality and stay time predictor for the eICU collaborative research database (Pollard et al., 2018) containing data of 208 hospitals and more than 200,000 patients.
-Although thousands of devices may be available in a given round, FL systems typically select a subset to participate in training.
-This selection process ensures a reasonable training time and accounts for diminishing returns where including more clients doesn’t accelerate convergence (Xu et al., 2024).
-Data heterogeneity is a major challenge in FL.
-Varying volumes and distribution of local data among participants is known to cause slow convergence and poor model accuracy (Yang et al., 2021; Wen et al., 2023).
-To mitigate the impact of heterogeneity, an effective approach is clustering, which groups statistically similar clients.
-Clients in the same cluster train one cluster model collaboratively and also have their inference requests served by this model.
-Previous works such as Auxo (Liu et al., 2023), FlexCFL (Duan et al., 2021), FedDrift (Jothimurugesan et al., 2023), and FedAC (Zhang et al., 2024) have shown that clustering can achieve high accuracy and fast model convergence.
-2.2. Data Drifts Affect Clustering
-The key challenges for clustered federated learning are data drifts.
-Data drifts refer to the changes in the percentage of clients’ local samples associated with a specific label.
-Drifts happen naturally when clients have access to non-stationary streaming data (e.g., cameras on vehicles and virtual keyboards on phones) (Koh et al., 2021; Bhardwaj et al., 2022).
-For example, when training an FL model for keyboard suggestions, drifts occur when many clients simultaneously start searching for a recent event (widespread drift), a few clients pick up a niche hobby (concentrated intense drift), or a new term appears when a product is launched (new label added).
-Data drifts make clustered FL ineffective as it increases the data heterogeneity in the clusters.
-We demonstrate this problem with the Functional Map of the World (FMoW) dataset (Christie et al., 2018).
-This dataset contains time-stamped satellite images labeled according to the function (e.g., recreational facility, crop field, etc.) of the captured land.
-We keep images taken on or after January 1, 2015, and create one client per unique UTM zone metadata value (302 clients in total).
-A day’s worth of data becomes available every two training rounds, and clients maintain images they received over the last 100 rounds.
-We use mean client distance as the intra-cluster heterogeneity metric (Lai et al., 2021).
-For each client, we calculate the average pairwise L1 distance between its data distribution vector and those of all other clients in the same cluster.
-We then use the mean of these per-client distances to quantify the overall heterogeneity after clustering.
-Note that instead of first averaging within each cluster and then finding the mean of those cluster averages, we use the overall mean across all clients.
-This helps us avoid biased results arising from imbalanced cluster sizes.
-Note that in the baseline case without any clustering, we consider all clients as members of a “global cluster”.
-Figure 1. Client heterogeneity of the global set and clusters generated through label-based clustering.
-We cluster clients based on their label distribution vectors at the starting round and track the overall heterogeneity by measuring the mean client distance as described earlier.
-We cluster based on label distribution because data drifts occurs when label distribution changes.
-Figure 1 shows that at the beginning (e.g., round 1), clustering reduces the per-cluster heterogeneity compared with no clustering (i.e., putting all clients in a global set).
-However, as data distribution shifts over time, static clustering increases per-cluster heterogeneity and soon gets close to the no clustering case at round 322.
-3. Fielding Design decisions
-Addressing data drifts is crucial for preserving the similarity of clients within a cluster.
-To handle data drifts, we must identify which clients should be moved, determine the appropriate clusters for these clients, and specify the information to collect for drift detection and client movements.
-3.1. Combine Per-client Adjustment and Re-clustering
-Existing clustered FL works rely on per-client adjustment or re-clustering to handle data drifts. However, neither approach works well on all levels of drifts.
-Per-client adjustment fails in many-drift scenarios.
-Per-client adjustment keeps a fixed number of clusters and moves drifted clients individually to another cluster with similar gradients or model weights (Ghosh et al., 2020; Duan et al., 2021).
-However, in highly dynamic environments where many clients change their distributions at the same time (Jothimurugesan et al., 2023), we have to move all these clients across clusters and the cluster membership changes dramatically.
-It is hard to decide which cluster the client should move to because the basis used by previous works to match client to the closest cluster (i.e., cluster’s average model weights or gradients) is changing (Ghosh et al., 2020; Duan et al., 2021).
-We demonstrate this problem in Figure 1 (“Move Individuals”).
-In rounds 1-40 and rounds 410-615, individual movements help keep the mean client distance below that of the global set even though clients are becoming more heterogeneous (signaled by the increasing heterogeneity value).
-However, as the global set of all clients experiences major drifts during rounds 600-730 (signaled by the heterogeneity spike), many clients start moving across clusters simultaneously.
-Per-client adjustment then results in poor clusters and an overall heterogeneity value higher than the global set in rounds 740-1340.
-Moreover, maintaining a fixed number of clusters can lead to sub-optimal clustering when new distribution patterns emerge during data drifts.
-For instance, we want to create another cluster when a new label appears for clients with a substantial number of samples of that label, but keeping the cluster number static will deem such creation infeasible.
-(a) Selective global re-clustering vs. default global re-clustering. Positive numbers mean selective re-clustering is better.
-(b) Global re-clustering vs. Selected re-clustering. Positive numbers mean re-clustering all drifted clients is better.
-Figure 2. Accuracy difference between different re-clustering approaches.
-Re-clustering makes model unstable during small drifts In Figure 1, we show an ideal case where we get optimal clusters by globally re-clustering all clients using label distribution vectors when any client drifts (“Global recluster”).
-Our implementation bounds the maximum number of clusters to 10, as previous works empirically show that having more clusters can negatively affect model convergence due to limited training resources (Liu et al., 2023).
-As expected, this approach yields the lowest heterogeneity value.
-However, we will next show that such low heterogeneity does not necessarily lead to superior model accuracy.
-Common clustering algorithms such as k-means have randomized initialization, leading to distinct results in multiple runs (Arthur and Vassilvitskii, 2006; Ahmed et al., 2020).
-This behavior is also observed in our experiment.
-In Figure 1, the “Global Recluster” curve exhibits more dramatic changes between rounds than other curves.
-Following the global re-clustering process, we need to determine the model each cluster should continue with.
-In our implementation, we align with previous works and use the average of the clients’ models within each cluster (Long et al., 2023; Zhang et al., 2024).
-When most clients remain in the same cluster, the new cluster models would be similar to the previous ones.
-But when there are significant changes in the clusters, either due to large drifts or bad centroid initialization, the new models will differ substantially and require additional time to stabilize.
-This leads to sudden accuracy drops as the new clusters warm up.
-In Figure 2(a), the x-axis denotes training rounds and the y-axis shows the mean accuracy we get with our selective global re-clustering (details in Section 3) subtracted by that of triggering global re-clustering by default.
-We observe that default global re-clustering leads to fluctuating test accuracy, which oscillates between meeting our approach’s accuracy and being as much as 5% worse.
-Summary.
-Per-client adjustment fails in highly dynamic environments where many clients experience drifts.
-In contrast, global re-clustering may harm model accuracy because it increases the instability of cluster memberships and increases the model convergence times, especially during minor drifts.
-When cluster memberships remain largely stable, global re-clustering is not only excessive but also requires additional steps to warm up the new clusters (e.g., finding optimal cluster model parameters).
-On the other hand, when many clients experience drifts, moving individual clients can cause substantial shifts in cluster centers and miss the opportunity to create or remove clusters as distribution patterns evolve.
-Based on the prior analysis, we opt for a design that starts with client individual migration and triggers global re-clustering selectively (Section 4 explains the detailed design).
-Our method is visualized as the red curve (“Move Ind.+Global Recluster”) in Figure 1.
-We retain a low heterogeneity value throughout training and achieve high model accuracy faster (Section  6).
-3.2. Re-clustering All Drifted Clients
-When we rerun the clustering algorithm every round, we can only easily recluster selected clients (Liu et al., 2023).
-Since selected clients already get the gradients or model parameters during training, it is easy to compare with the existing clusters and decide whether they experience drifts and need change.
-In contrast, since unselected clients do not report gradients, we cannot decide whether and where to move them.
-However, as mentioned in Section 2.1, pedagogical FL settings typically select only a subset of clients to contribute in each round.
-When unselected clients experience data drifts and are still stuck in their previous cluster, they might be served by the cluster model trained on data of a different distribution.
-We plot the difference in clients’ mean test accuracy of re-clustering only selected clients versus all clients that experience data drifts every round in Figure 2(b).
-Re-clustering all drifted clients offers a 1.5%-4% higher accuracy in most rounds, which is a significant difference given that the final accuracy we achieve on FMoW is 52.4%.
-This trend demonstrates that only moving selected clients leads to imperfect clusters containing diverging but unselected clients.
-Table 1.
-Heterogeneity of all clients and intra-cluster heterogeneity after gradient-based clustering (lower is better).
-At the initial stages, gradient-based clustering doesn’t sufficiently decrease heterogeneity due to the model being not well-trained.
-In contrast, label-based clustering provides a consistent decrease in heterogeneity.
-Total Rounds Un-Clustered Gradient-Based Clustering Label-Based Clustering
-100 1.81 1.82(+0.6%) 1.65(-8.8%)
-200 1.80 1.82(+1.1%) 1.64(-8.9%)
-500 1.78 1.71(-3.9%) 1.62(-9.0%)
-1000 1.74 1.66(-4.6%) 1.56(-10.3%)
-1500 1.76 1.63(-7.4%) 1.60(-9.1%)
-Global re-clustering based on training outputs incur high overheads.
-Considering the suboptimal accuracy of only re-clustering the selected clients, our goal is to move all drifted clients promptly.
-One effective method to learn information from all clients for re-clustering is to train a global model with selected clients and then distribute it to all clients.
-For example, in FlexCFL, each client performs local training with the global model weights on their own data and reports gradients to the central server, where clustering happens (Duan et al., 2021).
-The coordinator can then re-cluster all clients according to the gradients.
-This approach incurs high communication and computing overheads.
-For example, in our experiments training ResNet-18 (He et al., 2016) for image classification tasks, devices on average take 116.7 seconds to download the 11 million parameters and upload gradient coordinates.
-They also need an additional 50.4 seconds to run forward and backward propagation.
-3.3. Label-based Instead of Gradient-based Clustering
-Another issue with gradient-based re-clustering is that the clustering effectiveness is sensitive to the model quality.
-Table 1 presents the resulting mean client distance when we perform gradient-based clustering after training a global model for various numbers of rounds.
-The numbers in parentheses indicate the change in mean client distance relative to that of the global set.
-A negative value in the parentheses indicates that the generated clusters are overall less heterogeneous than the global set.
-When we perform gradient-based clustering with the global model at round 100 and round 200, the resulting clusters are more heterogeneous.
-As we postpone clustering to later rounds, gradient-based clustering achieves an increasingly larger reduction in mean client distance.
-This trend indicates that gradient-based clustering doesn’t perform well in earlier rounds, when the global model used to collect gradients from all clients hasn’t converged.
-We propose using label distribution vectors for clustering (we refer to this approach as “label-based”).
-Label distribution vectors are available on all clients and enable us to promptly detect any drift by checking for distribution changes.
-Compared to gradient-based clustering, label-based doesn’t demand parameter or gradient transmission, doesn’t introduce additional computation tasks, and is not sensitive to the timing of client clustering.
-In Table 1, label-based clustering provides consistent heterogeneity reduction across rounds.
-Furthermore, as a label distribution vector typically has tens or hundreds of coordinates compared to millions in a full gradient, label-based clustering incurs significantly lower computational overhead.
-In our experiment of training ResNet-18 on the FMoW dataset, clustering all 302 clients using our label-based solution takes only 0.05 seconds while the gradient-based solution FlexCFL takes 37.92 seconds (note that FlexCFL already accelerates this process through dimensionality reduction using truncated Singular Value Decomposition).
-4. Fielding Overview
-Figure 3.
-Clustering overview of Fielding.
-Clients send distribution vectors to the coordinator; the coordinator moves drifted clients to the closest cluster; global re-clustering is triggered if any cluster center shifts by a distance larger than $\tau$.
-To improve clustered FL performance, we develop Fielding, a clustered FL framework that efficiently reassigns clients after data drifts.
-As explained in Section 3, Fielding has three key design decisions: (1) reassign all drifted clients instead of only selected clients, (2) combine per-client adjustment and global re-clustering, and (3) cluster clients based on label distribution vectors.
-4.1. System Overview
-As discussed in Section 3.1, neither per-client adjustment nor global re-clustering works well across all drift scenarios.
-Therefore, we propose a combined approach that leverages both methods.
-Fielding employs a threshold $\tau$ to determine whether global re-clustering is necessary.
-When clients report data drifts, we first try re-clustering drifted clients individually by moving them to the cluster whose center is the closest to their new distribution vector.
-To ensure determinism across different orders of client movements, we do not update the cluster center throughout this per-client adjustment process.
-After individual movements, we recalculate the cluster centers and trigger global re-clustering over all clients only when any center moves by a distance larger than $\tau$.
-In our implementation, we set $\tau$ as one-third of the average distance between a pair of cluster centers, which is a value that works empirically well across evaluation settings.
-Figure 3 illustrates the clustering mechanism of Fielding.
-When clients register with the centralized coordinator, they send their local data distribution vector, either during initial registration or following a recent data drift.
-Upon receiving this vector, the coordinator updates the client’s metadata and assigns the client to the nearest cluster.
-Before selecting participants for the upcoming training round, the coordinator finds the distance each cluster center has shifted.
-It then initiates a global clustering process if any cluster exceeds the threshold $\tau$.
-Otherwise, it finalizes the updated cluster membership.
-Our clustering approach incurs additional communication costs, as clients are expected to upload their data distribution vector.
-However, this vector is small (4 bytes×$\times N$ where $N$ is the number of labels), and clients only need to send it when they first join or experience data drift.
-In Section 4.3, we prove the model convergence guarantees with label-based clustering.
-4.2. Algorithmic Framework
-1 Partition clients into clusters using Algorithm 3
+Fielding: Clustered Federated Learning with Data Drift
+Abstract
+Federated Learning (FL) trains deep models across edge devices without centralizing raw data, preserving user privacy.
+However, client heterogeneity slows down convergence and limits global model accuracy.
+Clustered FL (CFL) mitigates this by grouping clients with similar representations and training a separate model for each cluster.
+In practice, client data evolves over time – a phenomenon we refer to as data drift – which breaks cluster homogeneity and degrades performance.
+Data drift can take different forms depending on whether changes occur in the output values, the input features, or the relationship between them.
+We propose Fielding, a CFL framework for handling diverse types of data drift with low overhead.
+Fielding detects drift at individual clients and performs selective re-clustering to balance cluster quality and model performance, while remaining robust to malicious clients and varying levels of heterogeneity.
+Experiments show that Fielding improves final model accuracy by 1.9–5.9% and achieves target accuracy 1.16x–2.23x faster than existing state-of-the-art CFL methods.
+1 Introduction
+Federated Learning (FL) is a distributed machine learning paradigm that enables multiple clients to collaboratively train a shared model while keeping raw data local (Bonawitz et al., 2019; McMahan et al., 2017).
+FL has gained attention due to its privacy-preserving nature, as it avoids transmitting sensitive data to a central server (Lyu et al., 2024; Huba et al., 2022).
+This makes it particularly appealing in applications such as mobile text prediction (Yang et al., 2018), energy forecasting (Abdulla et al., 2024), and medical imaging (Zhou et al., 2023).
+In practice, FL involves clients with diverse hardware, usage patterns, and data sources, leading to variation in both dataset sizes and distributions: known as data heterogeneity (Li et al., 2022a; Kim et al., 2024).
+This heterogeneity slows model convergence and degrades model performance (Li et al., 2020a; Zhao et al., 2018; Sattler et al., 2020).
+Clustered FL (CFL) addresses this by grouping clients with similar data characteristics and training a separate model per cluster (Ghosh et al., 2020; Duan et al., 2021; Liu et al., 2023; Jothimurugesan et al., 2023), improving learning efficiency and accuracy by reducing intra-cluster heterogeneity.
+However, real-world deployments face data drifts over time.
+Data drift typically falls into three categories (Huyen, 2022): (1) label shift, where the output categories frequencies change but the characteristics associated with each category remain stable (Lipton et al., 2018; Garg et al., 2020); (2) covariate shift, where the distribution of the input features changes but the feature-label relationship remains constant (Ganguly and Aggarwal, 2023; Mallick et al., 2022); (3) concept shift, where the underlying input-output relationship changes, rendering a model trained on past data inaccurate on new data (Casado et al., 2022; Jothimurugesan et al., 2023).
+As data drift accumulates, intra-cluster heterogeneity increases, and clusters can become as diverse as the full client set (Jothimurugesan et al., 2023; Moreno-Torres et al., 2012).
+Effective drift handling must address two key challenges: adapting to varying drift magnitudes while maintaining practical efficiency, and supporting diverse drift types.
+The first challenge arises when drifts have varying magnitudes.
+When many clients experience drift, re-clustering based on outdated average weights or gradients becomes invalid, often requiring global re-clustering.
+FlexCFL (Duan et al., 2021) and IFCA (Ghosh et al., 2020) attempt to adapt incrementally by moving clients one at a time while keeping the number of clusters fixed.
+They, however, struggle under large-scale shifts as they reassign drifted clients using clusters’ current average gradient or weight, which shift significantly after massive reassignments.
+Global re-clustering reinitializes all clusters to reflect current data, but incurs significant communication and computation costs: clients might download multiple model replicas and upload new gradients; moreover, the newly formed clusters typically suffer an accuracy dip.
+Hence, while global re-clustering works for all drift magnitudes, incremental reassignment is more desirable under small drifts for practical reasons.
+FedDrift (Jothimurugesan et al., 2023) adopts global re-clustering and pays an amplified cost by having each client train on all cluster models.
+Auxo (Liu et al., 2023) and FedAC (Zhang et al., 2024) reduce overhead by only reassigning and adjusting cluster counts with clients selected for training, but ignore drifted clients that are not selected, reducing overall effectiveness (see Section 2.1).
+The second challenge is to detect and adapt to different types of drift.
+FL frameworks rely on client representations—client-side information used for clustering—to capture drifts.
+The label distribution vector is a common representation that reflects the label shift.
+Another is the input embedding: feature vectors derived from a client’s local data using a neural network, then aggregated (e.g., averaging) to represent the input distribution.
+When the input–output mapping remains stable, input feature shifts often correlate with label distribution shifts so that both representations can produce similar clustering results.
+However, neither captures changes in the input–output relationship itself – concept drift.
+Addressing concept drift typically requires loss-based representations, such as loss values, gradients, or gradient directions.
+For example, Sattler et al.
+(2021) clusters clients based on gradient directions to identify such shifts.
+Each representation has trade-offs: Label and embedding vectors capture only marginal distributions, while gradient-based features require additional computation, are sensitive to training noise, and evolve with model updates (Table 3 shows that gradient-based clustering generates better clusters as training progresses and the model becomes more stable).
+Selecting or combining these representations enables more reliable detection in all types of drift (see Appendix E).
+In this work, we present Fielding, a CFL framework that handles multiple types of data drift with low overhead.
+Fielding makes two key design choices: it combines per-client adjustment with selective global re-clustering, adapting to varying drift magnitudes without incurring the full cost of global re-clustering at every step; and it re-clusters all drifted clients using lightweight, drift-aware representations, keeping client-side computation and communication costs low while enabling efficient information collection from all clients.
+Our theoretical framework supports all representation choices and provides per-round utility guarantees along with convergence bounds, allowing Fielding to effectively handle all three types of data drift.
+To demonstrate practicality, we extended the FedScale engine (Lai et al., 2022) to support streaming data and built a Fielding prototype on top.
+We evaluated our framework on four image streaming traces with up to 5,078 clients experiencing various drift patterns.
+In these scenarios, Fielding improves the final accuracy by up to 5.9% and reaches target accuracy as much as 2.23× faster than standard CFL approaches.
+Moreover, we show that Fielding flexibly accommodates different drift types through pluggable client representations (e.g., label distribution vectors, input embeddings, and gradients), integrates seamlessly with a range of client selection and aggregation schemes, and remains robust against malicious clients and varying degrees of heterogeneity.
+2 Fielding
+Fielding is a CFL framework that handles multiple types of data drift with low overhead.
+Managing data drifts effectively requires identifying clients that need reassignment, determining their appropriate clusters, and collecting information to support drift detection and client movement.
+We begin by motivating the challenges of drift handling, followed by the design decisions that shape Fielding, and then provide an overview of the system.
+2.1 Motivation and design choices
+We analyze the limitations of existing clustering strategies under real-world drift to motivate the design of Fielding.
+Our analysis uses the Functional Map of the World (FMoW) dataset (Christie et al., 2018), containing 302 clients (one per unique UTM zone) with time-stamped satellite images labeled by land use (e.g., airport, crop field).
+Two training rounds correspond to one day.
+More details on FMoW appear in Section 3.
+To quantify intra-cluster heterogeneity, we use mean client distance (Lai et al., 2021): for each client, we compute the average pairwise L1 distance between its data distribution and that of clients in the same cluster.
+We then take the mean across clients.
+Figure 1: Client heterogeneity of the global set and label distribution-based clusters.
+(a) Selective vs. default global re-clustering (positive = selective better).
+(b) Global re-clustering vs. Selected clients re-clustering (positive = global better).
+Figure 2: Accuracy difference between different re-clustering approaches.
+Per-client adjustment fails under many drift.
+Per-client adjustment methods, such as IFCA (Ghosh et al., 2020) and FlexCFL (Duan et al., 2021), maintain a fixed number of clusters and reassign drifted clients individually based on gradients or model weights.
+As mentioned in Section 1, such assignments become unstable and lead to poor clustering when many clients drift simultaneously.
+Figure 1 (“Move Individuals”) highlights this behavior.
+Between rounds 600–730, a significant drift event increases intra-cluster heterogeneity, and from rounds 740–1340, per-client adjustment fails to restore effective clustering.
+In fact, heterogeneity exceeds that of the unclustered client set.
+Global re-clustering is unstable during small drifts.
+Global re-clustering resets all client assignments to reflect the current data, and it produces the lowest heterogeneity in our experiments (Figure 1, “Global Recluster”).
+However, due to the random initialization of k-means (Arthur and Vassilvitskii, 2006; Ahmed et al., 2020), the resulting clusters can vary considerably across rounds, even under small drifts.
+Combined with cluster warm-up, this leads to test accuracy fluctuations.
+Figure 2(a) shows the accuracy difference between our selective re-clustering approach (described in Section 2.2) and a baseline that triggers global re-clustering after every drift event.
+The baseline exhibits unstable accuracy, sometimes matching ours but falling short by up to 5%.
+Design choice: combine per-client adjustment with selective global re-clustering.
+Per-client adjustment breaks in highly dynamic environments where many clients experience drift.
+Meanwhile, global re-clustering can harm model accuracy by destabilizing cluster assignments and slowing convergence, particularly during minor drifts.
+To balance these trade-offs, Fielding combines both strategies: starting with individual client migration and triggering global re-clustering only when needed (see Section 2.2 for details).
+As shown by the red curve (“Move Ind.
++ Global Recluster”) in Figure 1, this approach maintains low intra-cluster heterogeneity throughout training (see Section 3).
+Clustering only selected clients reduces accuracy.
+Methods such as Auxo (Liu et al., 2023) and FedAC (Zhang et al., 2024) use training outputs—e.g., gradients or model parameters—from selected clients to re-cluster them without any additional cost.
+However, unselected clients do not provide such information, making it unclear whether they have drifted or where they should be reassigned to.
+As outlined in Appendix B.1, typical FL settings sample only a subset of clients in each round.
+When unselected clients drift but remain in their previous clusters, they may receive models trained on data with different distributions, reducing accuracy.
+Figure 2(b) compares the mean test accuracy when re-clustering only selected clients versus all drifted clients in each round.
+Re-clustering all drifted clients yields a 1.5%–4% accuracy gain across most rounds—a meaningful improvement given the final FMoW accuracy of 52.4%.
+These results show that re-clustering only selected clients produces imperfect clusters containing misaligned clients.
+Global re-clustering based on training outputs incurs high overheads.
+Handling all drifted clients requires input from every client, not just those selected for training.
+Depending on the representation, this method might incur significant overhead.
+In FlexCFL (Duan et al., 2021), clients train the global model locally and report gradients to the coordinator.
+In our experiments using ResNet-18 (He et al., 2016) on FMoW, clients spent on average 116.7 seconds for model download and gradient upload, and an additional 50.4 seconds running forward and backward passes on local data.
+Design choice: efficient re-clustering of all drifted clients across drift types.
+To handle drift comprehensively and efficiently, Fielding re-clusters all drifted clients, regardless of whether they were selected for training.
+Instead of relying on training outputs, it uses lightweight representations that can be collected from all clients with minimal overhead.
+To support different drift types while maintaining low cost, Fielding allows configurable client representations: it uses label-distribution vectors—small and easy to compute—for detecting label and covariate shifts, and relies on gradients only when needed to capture concept drift.
+This design avoids accuracy loss caused by stale cluster assignments and eliminates the cost of global model propagation and local retraining.
+2.2 System overview
+Fielding consists of two core components: a centralized coordinator that manages client representations, performs clustering, and orchestrates model training; and a client-side module that tracks local representations and reports updates when drift is detected.
+As discussed in Section 2.1, neither per-client adjustment nor global re-clustering alone performs well across all drift scenarios.
+Fielding adopts a hybrid approach: it uses per-client adjustment by default and triggers global re-clustering only when necessary based on a threshold $\tau$.
+When clients detect drift, they send updated representations to the coordinator.
+The coordinator initially reassigns each drifted client to the nearest cluster without updating cluster centers during this phase to ensure deterministic outcomes regardless of client processing order.
+After all individual adjustments, the coordinator recalculates cluster centers and measures the shift.
+If any center has moved more than $\tau$, global re-clustering is triggered for all clients; otherwise, the updated cluster membership is finalized.
+In our prototype, setting $\tau$ to one-third of the average inter-cluster distance worked empirically well.
+Figure 3 illustrates this clustering process using label-distribution vectors as representation.
+Clients submit their representations upon registration and whenever drift is detected.
+The coordinator updates client metadata and assigns each client to the nearest cluster.
+Before selecting participants for the next round, it checks whether any cluster centers have shifted beyond $\tau$ to determine if global re-clustering is needed.
+Figure 3: Clustering overview of Fielding with label distribution as client representation.
+Clients send distribution vectors to the coordinator; the coordinator moves drifted clients to the closest cluster and triggers global re-clustering if any cluster center shifts by a distance larger than $\tau$.
+2.3 Algorithmic framework
+In this section, we describe our algorithm – presented in Algorithm 1 – in detail.
+Initially, Fielding clusters the clients using the $k$-means algorithm.
+The distance between clients is measured based on clients representations; in particular, for label-based clustering, we measure distance between data distributions, which is computed as the $\ell_{1}$-distance between their histograms: that is, assuming we have $l$ labels, the distance between clients with label distribution $p_{1},\ldots,p_{l}$ and $q_{1},\ldots,q_{l}$ is defined as $\sum_{i=1}^{l}|p_{i}-q_{i}|$.
+We assume that the algorithm is executed for $T$ global iterations, and during each iteration, the label distribution of each client might change, for instance by adding or removing data points.
+The first step of Fielding is to handle data drift (Algorithm 2).
+Our goal is, by the end of each iteration, to maintain a good model for each cluster, with respect to its clients’ data.
+In this step, each drifted client is assigned to the closest cluster.
+If this causes a significant change in cluster centers – namely, if there exists a cluster moving by more than $\theta/3$, where $\theta$ is the average distance between cluster centers – we recluster clients from scratch.
+This condition avoids frequent reclustering, which might require excessive resources and can adversely change the clients’ losses.
+Importantly, we recluster all the clients, not just available clients, which, as we show in Appendix B, improves the accuracy.
+For the global reclustering, we choose the number of clusters that gives the highest silhouette score.
+After reclustering, we compute new cluster models: for each client $i$, let $\mathbf{x}_{i}$ be its old cluster model, and then for each new cluster $C_{k}$, we define its model as an average of $\mathbf{x}_{i}$ for $i\in C_{k}$ (i.e., we average the old client models).
+After the clusters are computed, we train the cluster models for $R$ rounds.
+In each round, we sample clients from the cluster and run $L$ local iterations of gradient descent.
+After local iterations, we set the cluster model as the average of the models of all participating clients.
+1 Partition clients into clusters using $k$-means clusters based on client representations
 2 Initialize cluster models $\mathbf{c}^{(1)},\ldots,\mathbf{c}^{(K)}$
 3 for every iteration $t=0,\ldots,T-1$ do
-4       Handle data drifts using Algorithm 2
+4       Handle data drift using Algorithm 2
 5       Let $C_{1},\ldots,C_{K}$ be the clusters
@@
 17      report cluster models and client-to-cluster assignment
-Algorithm 1 Clustered Federated Learning with data drift
+Algorithm 1 CFL with data drift
 1 for each drifting client $x$ do
@@
 7             Let $\mathbf{x}_{i}$ be the model corresponding to the $i$’th client’s cluster
-8      Let $C_{1},\ldots,C_{K}$ be clusters returned by Algorithm 3
+8      Let $C_{1},\ldots,C_{K}$ be $k$-means clusters based on client representations
 9       for each cluster $C_{k}$ do
@@
 11
-Algorithm 2 Handling data drifts
-1 Define the distance between clients as the $\ell_{1}$-distance between their label histograms
+Algorithm 2 Handling data drift
+1 Define the distance between clients based on their representations
 2 Choose $K$ with the largest silhouette score
@@
 Algorithm 3 Clustering
-In this section, we describe our algorithm – presented in Algorithm 1 – in detail.
-Initially, Fielding clusters the clients using the $k$-means algorithm (Algorithm 3).
-Importantly, we measure distance between data distributions, which is computed as the $\ell_{1}$-distance between their histograms: that is, assuming we have $l$ labels, the distance between clients with label distribution $p_{1},\ldots,p_{l}$ and $q_{1},\ldots,q_{l}$ is defined as $\sum_{i=1}^{l}|p_{i}-q_{i}|$.
-We assume that the algorithm is executed for $T$ global iterations, and during each iteration, the label distribution of each client might change, for instance by adding or removing data points.
-The first step of Fielding is to handle the data drifts (Algorithm 2).
-Our goal is, by the end of each iteration, to maintain a good model for each cluster, with respect to its clients’ data.
-In this step, each drifted client is assigned to the closest cluster.
-If this causes a significant change in cluster centers – namely, if there exists a cluster moving by more than $\tau/3$, where $\tau$ is the average distance between cluster centers – we use Algorithm 3 to recluster clients from scratch.
-This condition avoids frequent reclustering, which might require excessive resources and can adversely change the clients’ losses.
-Importantly, we recluster all the clients, not just available clients, which, as we show in Section 2, improves the accuracy.
-For the global reclustering, we choose the number of clusters that gives the highest silhouette score.
-After reclustering, we compute new cluster models: for each client $i$, let $\mathbf{x}_{i}$ be its old cluster model, and then for each new cluster $C_{k}$, we define its model as an average of $\mathbf{x}_{i}$ for $i\in C_{k}$.
-That is, we average the old client models.
-After the clusters are computed, we train the cluster models for $R$ rounds.
-During each round, we sample the clients from the cluster and run $L$ local iterations of gradient descent.
-After the local iterations, we set the cluster model as the average of the models of all participating clients.
-4.3. Convergence
-In Appendix A, we analyze the performance of our framework by bounding the average of clients’ loss functions at every iteration.
-Formally, for every cluster $C_{k}$, let $\mathbf{c}^{(k)}$ be the model corresponding to the cluster by the end of the iteration, and let $\mathbf{c}^{(k,*)}$ be the optimal model for this cluster, that is, the minimizer of $\sum_{i\in C_{k}}f^{(i)}_{t}$.
-Then, we bound the average loss of the clients with respect to the best possible cluster models:
-(1) $\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in C_{k}}(f^{(i)}_{t}(\mathbf{c}^{(k)})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}))$
-Our analysis has two ingredients, corresponding to two main stages in our algorithm: gradient descent steps and reclustering.
-For the gradient descent part, we use the standard convergence analysis of SGD, which for strongly convex function $f$ shows that for stochastic gradient descent update the following holds:
-$\mathbb{E}\left[f(\mathbf{x}_{T})-f(\mathbf{x}^{*})\right]$ $\leq\left(1-\eta\mu\right)^{T}(f(\mathbf{x}_{0})-f(\mathbf{x}^{*}))+\frac{L\eta}{2\mu}\sigma^{2}$
-The right-hand side has two terms, where the first term decays exponentially, and the second term corresponds to the total contribution of the stochastic noise.
-In our analysis, we similarly have the $(f(\mathbf{x}_{0})-f(\mathbf{x}^{*}))\left(1-\eta\mu\right)^{T}$ term which decays exponentially, and a non-vanishing term corresponding to the noise coming from the stochastic noise and the client sampling.
-The major challenge in the analysis is bounding the effect of data drifts and reclustering.
+2.4 Convergence
+In Appendix A, we analyze the performance of simplified version of our framework by bounding the average of clients’ loss functions at every iteration.
+The major challenge in the analysis is bounding the adverse effects of data drifts and reclustering.
 Data drifts change client datasets, hence changing their local objective functions.
-On the other hand, while reclustering is useful in the long run, its immediate effect on the objective (1) can potentially be negative (see Figure 2(b)) due to the mixing of models from different clusters.
-To address these issues, we introduce natural assumptions.
-First, we assume that the distance between client representations – which in our algorithm correspond to label distributions – translates into the difference between their objective functions.
+On the other hand, while reclustering is useful in the long run, its immediate effect on the objective can potentially be negative (see Figure 2(b)) due to the mixing of models from different clusters.
+We make the following assumptions.
+[^1] See Appendix A for the precise statements and the discussion of the assumptions First, we assume that the distance between client representations – for example label distributions – translates into the difference between their objective functions.
 Second, we assume that the effect of data drift is bounded: that is, the representation $r_{i}$ of client $i$ changes by at most $\delta$ for each data drift.
-Finally, we assume that the clients are clusterable; that is, $K$ clusters exist, so representations of clients within each cluster are similar.
+Finally, we assume that clients are clusterable; that is, there exists $K$ clusters so that representations of clients within each cluster are similar.
 Assumption A.
-$f^{(i)}_{t}$ – local objective for client $i$ at iteration $t$ – is $\mu$-strongly convex and $L$-smooth, and we have access to a stochastic oracle with variance $\sigma^{2}$.
-There exists clustering such that representations inside each cluster are $\Delta$-close, the data drift changes representations by at most $\delta$, and the difference between the objective functions is at most the distance between representations multiplied by $\theta$.
-Under these assumptions, the data drift can affect each client representation – and hence objective 1 – by at most a fixed value.
-Moreover, by the clustering assumption, the clients within each cluster have similar objective function; moreover, this holds both before and after the reclustering.
-This allows us to bound the increase in the loss function due to reclustering.
-In Appendix A, we prove the following theorem.
-Theorem 4.1.
+Each $f^{(i)}_{t}$ – the local objective for client $i$ at iteration $t$ – is $L$-smooth and satisfies $\mu$-Polyak-Łojasiewicz condition, and we have access to a stochastic oracle with variance $\sigma^{2}$.
+There exists clustering such that representations inside each cluster are $\Delta$-close, the data drift changes representations by at most $\delta$, and the ratio between the difference between the objective function values and the distance between representations is at most $\theta$.
+Under these assumptions, data drifts can affect each client representation – and hence the objective we optimize – by at most a fixed value.
+Moreover, by the clustering assumption, clients within each cluster have similar objectives after the reclustering, allowing us to bound the increase in the loss function due to reclustering.
+In Appendix A, we show the following result.
+Theorem 1.
 Let $N$ be the number of clients and $M$ be the total number of machines sampled per round.
@@
 Then, under A, for $\eta\leq 1/L$, for any iteration $t$ we have
-$\frac{1}{N}\sum_{k\in[K]}\sum_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t+1})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$
-$\leq(1-\eta\mu)^{TR}\left(f_{0}(\mathbf{x}_{0})-f_{0}(\mathbf{x}^{*})\right)$
-$+\frac{L\eta}{2\mu}\left(\frac{\sigma^{2}+8L\theta\Delta}{M/K}+3\theta(\Delta+\delta)(1-\eta\mu)^{R}\right)$
-That is, objective (1) has two terms: an exponentially decaying term corresponding to the initial loss, and a non-vanishing term corresponding to stochastic noise, as well as loss due to clustering and data drift.
-5. Implementation
-We implemented Fielding in Python on top of FedScale (Lai et al., 2022), a state-of-the-art open-source FL engine.
-We follow FedScale’s design of having one centralized coordinator and client-specific executors.
-Figure 4 illustrates the end-to-end workflow of Fielding.
-At the beginning of each training round, 1 client executors register with the coordinator to participate and report their local data distribution optionally.
-The coordinator’s cluster manager maintains clients’ distribution records and 2 moves clients to the closest cluster when their reported distribution drifts.
-After moving drifted clients individually, the cluster manager measures center shift distances and either triggers global re-clustering when any cluster center shifts significantly or finalizes clusters membership.
-In the case of global clustering, we use $K$-means clustering and determine $K$ using the silhouette method (Rousseeuw, 1987).
-To stabilize new clusters and avoid creating poor clusters due to bad centroids initialization (Ahmed et al., 2020), we reuse previous cluster centers when the target $K$ value is no greater than the number of existing clusters.
-The initial model of each newly created cluster is set as the average of its clients’ previous cluster model (see Algorithm 2).
-Figure 4.
-The system architecture and workflow of Fielding.
-Blue arrows represent communication between clients and the coordinator.
-Black arrows indicate the workflow among different components on the same machine.
-When client clustering is done, the coordinator notifies the client selector to 3 select a subset of clients to contribute to each cluster and 4 communicate the latest model parameters to the selected clients.
-5 Upon receiving a set of parameters, client executors conduct the training process over local data and 6 upload the updated model parameters to the coordinator.
-Finally, 7 the model manager aggregates individual models and stores the new cluster models.
-This process happens iteratively until the clients’ mean test accuracy reaches a target value.
-Fielding also regularly creates checkpoints for the models, clients’ metadata, and cluster memberships for future fine-tuning and failure recovery.
-Fielding’s overheads are mainly determined by the number and size of distribution vectors we need to re-cluster.
-In our largest evaluation setting where we train with 5078 clients on a dataset with 100 labels, per-client adjustment takes 2.0 seconds and global re-clustering takes 15.6 seconds on average.
-Storing the latest reported distribution vector for each client consumes $5078\times 100\times 4$B $\approx 1.9$MB of memory.
-6. Evaluation
-We evaluate Fielding’s effectiveness by running four FL tasks.
-We also demonstrate Fielding’s compatibility with various client selection strategies and aggregation algorithms.
-Our end-to-end results highlight that
-  • Fielding improves the final test accuracy by 1.9%-5.9% and is more stable than other clustered FL works when facing real-world data drifts.
-  • Fielding works well with complementary FL optimizations and offers consistent accuracy improvements across different settings.
-  • Fielding is robust to malicious clients who report corrupted distribution vectors and accelerates model convergence when clients are less heterogeneous.
-Figure 5. Time to accuracy (TTA) comparison over four tasks.
+$\frac{1}{N}\sum_{k\in[K]}\sum_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t+1})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$ $\leq(1-\eta\mu)^{tR}\left(f_{0}(\mathbf{x}_{0})-f_{0}(\mathbf{x}^{*})\right)$
+$\quad+\frac{L\eta}{2\mu}\left(\frac{\sigma^{2}+8L\theta\Delta}{M/K}+3\theta(\Delta+\delta)(1-\eta\mu)^{R}\right)$
+Intuitively, at every iteration, we provide a “regret” bound, comparing the loss at each cluster with the best loss we could have at each cluster.
+The bound has two terms: an exponentially decaying term corresponding to the initial loss, and a non-vanishing term corresponding to stochastic noise and the loss due to clustering and data drift.
+Ultimately, at early stages, the first term dominates, and we observe rapid improvement in objective value.
+On the other hand, at later stages, the first term vanishes, and the behavior is dictated by the stochastic and sampling noise, and severity of data drifts.
+3 Evaluation
+We evaluate Fielding prototype on four FL tasks with label distribution vectors as client representations (see Appendix E for results with other representations).
+End-to-end results highlight that Fielding improves final test accuracy by 1.9%-5.9% and is more stable than prior CFL methods.
+It works well with complementary FL optimizations, remains robust against malicious clients reporting corrupted distribution vectors, and still accelerates model convergence under low heterogeneity.
 Environment.
-We emulate large-scale FL training with four NVIDIA A100 GPUs.
-We use FedScale’s device datasets which provide realistic device computing and network capacity numbers while following the standardized training configurations in the original paper (Lai et al., 2022).
-Per-device computation time estimation is based on device computing speed, batch size, and number of local iterations.
-And we estimate the communication time as the volume of downloaded and uploaded data divided by the network bandwidth.
+We emulate large-scale FL training with two GPU servers, each with two NVIDIA A100 GPUs (80 GB memory) and two AMD EPYC 7313 16-core CPUs.
+We use FedScale’s device datasets for realistic device computing and network capacity profiles while following the standardized training setup in the original paper (Lai et al., 2022).
+Per-device computation time is estimated from device computing speed, batch size, and number of local iterations; communication time is estimated based on the volume of downloaded and uploaded data and network bandwidth.
 Datasets and models.
-We use a satellite image dataset (FMoW (Christie et al., 2018)) and two video datasets (Waymo Open (Sun et al., 2020) and Cityscapes (Cordts et al., 2016)) that include natural data drifts.
-For FMoW, we keep images taken after January 1, 2015, and create one client for each UTM zone.
-Two training rounds correspond to one day in this dataset.
+We use a satellite image dataset (FMoW (Christie et al., 2018)) and two video datasets (Waymo Open (Sun et al., 2020) and Cityscapes (Cordts et al., 2016)) that exhibit natural data drift.
+For FMoW, we keep images taken after January 1, 2015 and create one client per UTM zone.
+Two training rounds correspond to one day.
 For Waymo Open and Cityscapes, we use video segments pre-processed by Ekya (Bhardwaj et al., 2022).
 We create one client per camera per Waymo Open segment (212 clients in total), and one client per 100 consecutive frames per Cityscapes segment (217 clients in total).
-As these two datasets have video frame IDs within segments but lack global timestamps, we sort the samples on each client by frame ID and then partition the frames into 10 intervals.
-We stream in one data interval every 20 training rounds.
-We train ResNet-18 (He et al., 2016) on these three real-world datasets.
-To demonstrate how Fielding behaves in highly dynamic settings, we construct an artificial trace using the FedScale Open Images (Kuznetsova et al., 2020) benchmark.
-We find the top 100 classes in terms of number of samples and keep clients whose local data labels include at least 10 of the top 100 classes (5078 clients in total).
-Each client randomly and independently partitions all local data labels into 10 buckets.
-Each bucket then contains all local samples of labels mapped to it.
-We stream in one bucket every 50 training rounds.
-We train ShuffleNet v2 (Ma et al., 2018) on Open Images.
-On all four datasets, clients retain samples they received over the last 100 rounds.
-The details of experiment settings and training parameters are included in Appendix B.
+As these two datasets have video frame IDs within segments but lack global timestamps, we sort samples by frame ID and split them into 10 intervals.
+We stream in one data interval every 30 rounds with Cityscapes and every 20 rounds with Waymo Open.
+We train ResNet-18 (He et al., 2016) on Cityscapes and FMoW, and VisionTransformer-B16 (Dosovitskiy et al., 2021) on Waymo Open.
+To evaluate Fielding under highly dynamic drift, we construct a synthetic trace using the FedScale Open Images (Kuznetsova et al., 2020) benchmark.
+We find the top 100 most frequent classes and retain clients with samples from at least 10 of these 100 classes (5078 clients in total).
+Each client randomly partitions local data labels into 10 buckets, with each bucket then containing all samples of its labels.
+We stream in one bucket every 50 rounds and use ShuffleNet v2 (Ma et al., 2018) for this task.
+For all four datasets, clients start with 100 rounds worth of data and retain samples from the most recent 100 rounds.
+Dataset and training configurations are summarized in Table 1 and Table 2.
 Baselines.
@@
 Our main evaluation metrics are Time-to-Accuracy (TTA) and final test accuracy.
-We define TTA as the training time required to achieve the target accuracy, which is the average test accuracy when training a single global model for all clients.
+We define TTA as the training time required to achieve the target accuracy–the average client test accuracy when training a single global model (black dashed lines in Figure 4).
 We report the final average test accuracy across diverse settings to demonstrate the sensitivity and robustness of Fielding.
-6.1. End-to-end Training Performance
-Figure 5 shows that Fielding improves clients’ average test accuracy by 5.9%, 1.9%, and 2.1% on FMoW, Cityscapes, and Waymo Open respectively.
-When considering the global model’s final test accuracy as the target accuracy, Fielding offers a 1.27$\times$, 1.16$\times$, and 2.61$\times$ training speed up respectively (we define the moment when Fielding’s accuracy starts invariably exceeding the target accuracy as the point at which it achieves the target accuracy).
-On the highly heterogeneous and dynamic Open Images trace, Fielding achieves a significant accuracy improvement of 26%.
-Although Auxo exhibits comparable or better performance than the global model baseline, the benefits are limited: accuracy improvement up to 0.9% and convergence speedup up to 1.05$\times$.
-These results are less impressive than those reported in the original paper on static datasets, suggesting a decline in clustering effectiveness.
-This degradation stems from re-clustering only the selected participants after each round, which neglects clients that have drifted but remain unselected.
-Additionally, this approach can lead to sudden accuracy drops on the biased Waymo Open dataset, where over 80% samples are cars, as selecting diverging clients can cause dramatic model weight changes.
-FlexCFL fails to improve model accuracy or convergence speed on all three real-world datasets.
-This is because FlexCFL migrates clients individually without adjusting the number of clusters.
-As we observe in Figure 1, such an approach will eventually make clusters as heterogeneous as a global set containing all clients.
-(a) Oort selection
-(b) Distance-based selection
-Figure 6. Fielding with different client selection strategies on FMoW dataset.
-6.2. Compatibility with Other Optimizations
+Figure 4: Time to accuracy (TTA) comparison over four tasks.
+3.1 End-to-end training performance
+Figure 4 shows that Fielding improves average client test accuracy by 5.9%, 1.9%, and 3.4% in FMoW, Cityscapes, and Waymo Open, respectively.
+These improvements are significant, given that baseline test accuracies range from 45% to 85%, and are comparable to the accuracy gains prior CFL works achieve on static data.
+When considering the global model’s final test accuracy as the target accuracy, Fielding offers a 1.27$\times$, 1.16$\times$, and 2.23$\times$ training speed-up, respectively (We define the point at which Fielding’s accuracy consistently surpasses the target accuracy as the moment it achieves the target accuracy).
+On the highly heterogeneous and dynamic Open Images trace, Fielding boosts accuracy by 17.9-26.1% in the final 100 rounds.
+While Auxo performs comparably or slightly better than the global model baseline, the benefits are limited: up to 0.9% accuracy gain and up to 1.05$\times$ convergence speedup.
+These results fall short of those reported in the original paper on static datasets, suggesting a decline in clustering effectiveness.
+This degradation stems from re-clustering only the selected participants after each round, ignoring drifted but unselected clients.
+This approach also leads to sudden accuracy drops on the biased Waymo Open dataset, where over 80% samples are cars, due to sudden model shifts when divergent clients are selected.
+FlexCFL shows no improvement in accuracy or convergence across all real-world datasets, as it migrates clients individually without adjusting the number of clusters.
+As Figure 1 shows, this approach leads to cluster heterogeneity approaching that of the global client set.
+FedDrift Comparison.
+FedDrift (Jothimurugesan et al., 2023) is another clustered FL method designed to handle drift.
+However, it is impractical for large-scale settings, as it (1) requires every client to train every cluster model in each round and (2) assumes all clients remain online and available for training.
+Therefore, we construct a small-scale setting with the FMoW dataset and 50 clients (the number of clients we selected per round in our original experiment).
+Figure 10 shows that Fielding achieves 2.5% higher final accuracy and reaches FedDrift’s final accuracy 2.01$\times$ faster.
+3.2 Compatibility with other optimizations
 To demonstrate that Fielding is agnostic to client selection and model aggregation strategies, we evaluate Fielding’s performance on FMoW together with various complementary optimizations.
-In Figure 6 and Figure 7, we use the algorithm name alone to denote the baseline where we train one global model and use “+ Fielding ” to denote running the algorithm atop Fielding.
+In Figure 5 and Figure 6, we use the algorithm name alone to denote the baseline where we train one global model and use "+ Fielding " to denote running the algorithm atop Fielding.
 We use the Oort selection algorithm (Lai et al., 2021) and a distance-based algorithm prioritizing clients closer to the distribution center as client selection examples.
-As shown in Figure 6, Fielding improves the final average test accuracy by 6.0% and 4.5% while reaching the target accuracy 2.62×2.62$\times$2.62 × and 1.17×1.17$\times$1.17 × faster.
+As shown in Figure 5, Fielding improves the final average test accuracy by 6.0% and 4.5% and reach the target accuracy 2.62×2.62$\times$2.62 × and 1.17×1.17$\times$1.17 × faster.
 Note that Oort selection actively incorporates clients otherwise filtered out due to long response time, leading to a longer average round time.
-We also show that Fielding works well with aggregation algorithm FedYogi (Reddi et al., 2021) and q-FedAvg (Li et al., 2020c) in Figure 7.
+Figure 6 shows that Fielding works well with aggregation algorithm FedYogi (Reddi et al., 2021) and q-FedAvg (Li et al., 2020c).
 Fielding improves the final accuracy by 5.9% and 4.9% while giving a 1.34×1.34$\times$1.34 × and 1.25×1.25$\times$1.25 × speedup respectively.
-6.3. Robustness and Sensitivity Analysis
+3.3 Robustness and sensitivity analysis
 Malicious clients.
@@
 As reported in Figure 10, Fielding consistently outperforms the baseline global model accuracy across different percentages of malicious clients.
+(a) Oort selection
+(b) Distance-based selection
+Figure 5: Fielding with different client selection strategies on FMoW dataset.
 (a) FedYogi
 (b) q-FedAvg
-Figure 7. Fielding with different FL algorithms on FMoW dataset.
-Figure 8. Fielding is robust to malicious clients.
-Figure 9. Fielding offers gains across heterogeneity degrees.
-Figure 10. Fielding offers gains across heterogeneity degrees.
+Figure 6: Fielding with different FL algorithms on FMoW dataset.
+Figure 7: Small-scale comparison against FedDrift on FMoW.
+Figure 8: Fielding is robust to malicious clients.
+Figure 9: Fielding offers gains across heterogeneity degrees.
+Figure 10: Fielding improves model accuracy on static data.
 Varying data heterogeneity.
@@
 This is likely due to Auxo re-clustering all selected clients after each round by default, causing unnecessary cluster membership changes when the data is completely static.
-7. Related Work and Discussion
-Clustered FL Previous clustering solutions (Ghosh et al., 2020; Duan et al., 2021; Liu et al., 2023; Zhang et al., 2024; Jothimurugesan et al., 2023) have achieved success at bounding intra-cluster heterogeneity and improving model performance.
-However, their improvements are compromised by real-world data drifts as they either assume stable environments or handle drifts insufficiently.
+4 Related works
 Heterogeneity-aware FL.
-Recent works have tackled heterogeneity challenges through client selection, workload scheduling, and update corrections.
+Recent works have also tackled heterogeneity challenges through client selection, workload scheduling, and update corrections.
 Oort considers both statistical and system utility and designs an exploration-exploitation strategy for client selection (Lai et al., 2021).
@@
 Drifts-aware FL.
-Adaptive-FedAvg is an FL algorithm that adapts client learning rate based on the model variability between consecutive rounds (Canonaco et al., 2021).
-CDA-FedAvg detects concept drifts and extends FedAvg for learning under such drifts.
-It defines a short-term and a long-term memory for each client and applies rehearsal using data in the lone-term memory when drifts happen (Casado et al., 2022).
-Master-FL proposes a multi-scale algorithmic framework that instructs clients to train over multiple time horizons with a suitable learning rate for each time chunk (Ganguly and Aggarwal, 2023).
-Fielding focuses on dynamic client clustering that is robust against potential data drifts.
+Adaptive-FedAvg handles concept drift by extending FedAvg with a learning rate scheduler that considers the variance of the aggregated model between two consecutive rounds (Canonaco et al., 2021).
+CDA-FedAvg detects concept drifts and extends FedAvg with a short-term and a long-term memory for each client.
+It applies rehearsal using data in the lone-term memory when drifts happen (Casado et al., 2022).
+Master-FL proposes a multi-scale algorithmic framework that trains clients across multiple time horizons with adaptive learning rates (Ganguly and Aggarwal, 2023).
+They are complementary to our work and can improve cluster models when minor drifts that don’t trigger global re-clustering happen.
+Fielding focuses on dynamic client clustering that is robust against potential data drift.
 It is agnostic to other FL optimizations and should work seamlessly with the FL aggregation, selection, and scheduling algorithms above.
-8. Conclusion
-We have presented Fielding, a novel clustering-based Federated Learning framework designed to handle data drifts with low overhead.
-Fielding addresses the challenges of data drift by re-clustering clients based on their label distribution vectors as their data distributions change, dynamically adjusting the number of clusters, and stabilizing responses to minor data shifts.
-Fielding is robust against malicious clients and varying degrees of heterogeneity, and it is compatible with various client selection and aggregation algorithms.
-In Fielding, clients register with a centralized coordinator that updates their metadata and assigns them to the nearest cluster based on their local data distribution.
-Before each training round, the coordinator determines how much each cluster center has shifted.
-If any cluster exceeds a predefined threshold, a global re-clustering process is initiated; otherwise, the updated cluster memberships are finalized.
-Experimental results demonstrate that Fielding improves the final test accuracy by 1.9% to 5.9% compared to existing methods and provides greater stability when facing real-world data drifts across different settings.
-There are potential directions to further enhance Fielding.
-Our current approach requires labeled data, which may not be available in all contexts.
-Additionally, clients actively report their label distribution vectors to the coordinator, which introduces latent privacy concerns.
-In future work, we will focus on alternatives, such as inferring distribution vectors from gradients.
+Clustered FL.
+FlexCFL (Duan et al., 2021) and IFCA (Ghosh et al., 2020) adjust clusters by relocating shifted clients while maintaining a fixed number of clusters, operating under the assumption of low drift magnitude.
+Auxo (Liu et al., 2023) and FedAC (Zhang et al., 2024) re-cluster selected clients using gradient information or local models, delaying drift adjustments for unselected clients.
+FedDrift (Jothimurugesan et al., 2023) re-clusters all clients by broadcasting all cluster models every round to compute client local training loss, leading to significant costs.
+Appendix B discusses other prior CFL works without drift handling and the impact of drifts on static clusters.
+5 Conclusion
+We presented Fielding, a clustering-based Federated Learning framework designed to handle multiple types of data drift with low overhead.
+Effective drift handling requires adapting to varying drift magnitudes while maintaining efficiency and supporting diverse drift types.
+To meet these goals, Fielding makes two key design choices: it combines per-client adjustment with selective global re-clustering and re-clusters all drifted clients using lightweight, drift-aware representations.
+Fielding dynamically adjusts the number of clusters and stabilizes the system’s response to minor drifts.
+It is robust to malicious clients and varying levels of heterogeneity, and remains compatible with a range of client selection and aggregation strategies.
+In Fielding, clients register with a centralized coordinator, which maintains metadata and assigns each client to the nearest cluster based on its local representation.
+Before each training round, the coordinator checks for shifts in cluster centers.
+If any center moves beyond a predefined threshold, global re-clustering is triggered; otherwise, the system proceeds with updated cluster assignments, significantly reducing the overall re-clustering cost and mitigating loss increase after re-clustering.
+Experimental results show that Fielding improves final test accuracy by 1.9%–5.9% over existing methods and provides greater stability under real-world data drift.
+Limitations.
+There are potential directions to further enhance Fielding from the perspective of client representation.
+Reporting these representations – label distributions, embeddings, or gradients – poses potential privacy concerns.
+The differentially private approaches for private clustering are hard to apply in our settings since we perform clustering multiple times.
+Additionally, gradient representation, while being able to capture concept drifts, leads to significant overhead.
+We will address these concerns in future work.
 Appendix A Proof of convergence
-1 input: the number of clients $K$, the number of data drift events $T$, the number of sampled clients per round $M$, the number of rounds per data drift event $R$
+1 input: the number of clusters $K$, the number of data drift events $T$, the number of sampled clients per round $M$, the number of rounds per data drift event $R$, reclustering threshold $\Delta$
 2 Initialize cluster models $\mathbf{c}^{(1)}_{0},\ldots,\mathbf{c}^{(K)}_{0}$ with the same model $\mathbf{x}_{0}$
@@
 7             Let $\mathbf{x}_{i}$ be the model corresponding to the cluster containing client $i$
-8      Let $C_{1},\ldots,C_{K}$ be the $K$-center clustering of all clients based on representations from model $\tilde{\mathbf{x}}_{\tau}$
+8      Let $r^{(1)}_{t},r^{(2)}_{t},\ldots$ be the client representations
+9       if there exists a cluster $C$ such that $\rho(r^{(i)}_{t},r^{(j)}_{t})>\Delta$ for clients $i,j\in C$ then
+10             Let $C_{1},\ldots,C_{K}$ be the $K$-center clustering of all clients based on representations
+11      else
+12             Let $C_{1},\ldots,C_{K}$ be the current clusters
 13      for $k=1,\ldots,K$ do
-10             $\tilde{\mathbf{c}}^{(k)}_{0}\leftarrow\operatorname*{avg}_{i\in C_{k}}\mathbf{x}_{i}$
+14             $\tilde{\mathbf{c}}^{(k)}_{0}\leftarrow\operatorname*{avg}_{i\in C_{k}}\mathbf{x}_{i}$ // Cluster model is the average of clients’ models
 15      for $\tau=0,\ldots,R-1$ do
@@
 18                   $\tilde{\mathbf{c}}^{(k)}_{\tau+1}\leftarrow\tilde{\mathbf{c}}^{(k)}_{\tau}-\eta\operatorname*{avg}_{i\in S}G^{(i)}_{t}(\tilde{\mathbf{c}}^{(k)}_{\tau})$
-15
+19
 20      for $k=1,\ldots,K$ do
 21             $\mathbf{c}^{(k)}_{t+1}\leftarrow\tilde{\mathbf{c}}^{(k)}_{R}$
-18
+22
 Algorithm 4 Clustered Federated Learning with Data Drift
@@
 (2020a).
-In the algorithm, we recluster clients after every data drift event and use the $k$-center clustering instead of $k$-means clustering to simplify the proof.
+To simplify the proof, in the algorithm, we make the following changes compared with the algorithm from the main body: 1) we use the $k$-center clustering instead of $k$-means clustering; 2) we perform reclustering when the intra-clsuter heterogeneity exceeds a certain threshold.
+We compare this threshold condition with the one from the main body in Section F.2.
 We present the algorithm in Algorithm 4.
 After every data drift event, we first accept data changes for each client.
-After that, we recluster the clients, and assign to each cluster a model by averaging models of all clients in the cluster.
+After that, if the heterogeneity within some clsuter exceeds a certain threshold, we recluster the clients, and assign to each cluster a model by averaging models of all clients in the cluster.
 After that, for several rounds, we perform standard federated averaging updates on each cluster: we sample $M$ clients, compute the stochastic gradient for each client, and update the cluster model with the sampled gradients.
 Assumption B (Objective functions).
-Let $f^{(i)}_{t}$ be the local function at each client $i$ at data drift event $t$. Then $f^{(i)}_{t}$ is $\mu$-strongly convex and $L$-smooth (has an $L$-Lipschitz gradient) for all $i$ and $t$.
-The following are the standard assumptions on stochastic gradient.
+Let $f^{(i)}_{t}$ be the local function at each client $i$ at data drift event $t$. Then, for all $i$ and $t$:
+  • $f^{(i)}_{t}$ is $L$-smooth: $\|\nabla f^{(i)}_{t}(\mathbf{x})-\nabla f^{(i)}_{t}(\mathbf{x})\|\leq L\left\|\mathbf{x}-\mathbf{y}\right\|$ for all $\mathbf{x},\mathbf{y}$
+  • $f^{(i)}_{t}$ satisfies $\mu$-Polyak-Łojasiewicz ($\mu$-PL) condition: $\|f^{(i)}_{t}(\mathbf{x})\|^{2}\geq 2\mu(f^{(i)}_{t}(\mathbf{x})-f^{(i)}_{t}(\mathbf{x}^{*}))$, where $\mathbf{x}^{*}$ is the minimizer of $f^{(i)}_{t}$.
+The $\mu$-PL condition is a relaxation of the strong convexity assumption; both asumptions are standard in the federated learning literature (Cho et al., 7 29; Li et al., 2020a; Cho et al., 2020).
+The following are standard assumptions on stochastic gradient.
 Assumption C (Stochastic gradients).
-For each client $i$, at every data drift event $t$, we have access to stochastic gradient oracle $G^{(i)}_{t}$:
-  • $\mathbb{E}\left[G^{(i)}_{t}(\mathbf{x})\right]=f^{(i)}_{t}(\mathbf{x})$ for all $\mathbf{x}$;
-  • $\mathbb{E}\left\|G^{(i)}_{t}(\mathbf{x})-f^{(i)}_{t}(\mathbf{x})\right\|^{2}\leq\sigma^{2}$ for all $\mathbf{x}$.
-The next assumption connects the objective value with client representations.
-Client representations can take various forms, and in our paper, we assume that the representation of the cluster is its label distribution.
+For each client $i$, at every data drift event $t$, let $G^{(i)}_{t}$ be the stochastic gradient oracle:
+  • $\mathbb{E}\left[G^{(i)}_{t}(\mathbf{x})\right]=\nabla f^{(i)}_{t}(\mathbf{x})$ for all $\mathbf{x}$;
+  • $\mathbb{E}\left\|G^{(i)}_{t}(\mathbf{x})-\nabla f^{(i)}_{t}(\mathbf{x})\right\|^{2}\leq\sigma^{2}$ for all $\mathbf{x}$.
+The next assumption connects the objective value with client representations. Client representations can take various forms, such as label distributions, input embeddings, or gradients.
 Assumption D (Representation).
 For a metric space $(X,\rho)$, let $r^{(i)}_{t}\in X$ be a representation of client $i$ at time $t$.
-Then there exists a constant $\theta$ such that $\left|f^{(i)}_{t}(\mathbf{x})-f^{(j)}_{t}(\mathbf{x})\right|\leq\theta\cdot\rho\left(r^{(i)}_{t},r^{(j)}_{t}\right)$ for all $i,j,t$.
+Then there exists $\theta$ such that $\left|f^{(i)}_{t}(\mathbf{x})-f^{(j)}_{t}(\mathbf{x})\right|\leq\theta\cdot\rho\left(r^{(i)}_{t},r^{(j)}_{t}\right)$ for all $i,j,t$.
+We next give the intuition of why the assumption is natural for the label-based and the embedding-based distnaces.
+If the concept $P(\text{output}|\text{input})$ doesn’t change between clients, every model achieves the same expected error on each class.
+For each class, assuming that the class data for different clients is sampled from the same distribution, we get the same expected loss on this class for different clients.
+Hence, the overall difference in the loss function is largely determined by the fraction of each class in the clients’ data.
+Similarly, if the client embedding faithfully captures information about its inputs, assuming the trained models depend on the inputs continuously, they have the same expected error on similar inputs.
 Clearly, if data on clients can change arbitrarily at every data drift event, in general, there is no benefit in reusing previous iterates.
@@
 At every data drift event, the representation of each client changes by at most $\delta$: $\rho\left(r^{(i)}_{t},r^{(i)}_{t+1}\right)\leq\delta$ for all $i$, $t$.
+This assumption naturally holds in case when, for each client, only a small fraction of data points changes at every round.
 Finally, we assume that the clients are clusterable: there exist $K$ clusters so that within each cluster all points have similar representation.
@@
 At every data drift event, the clients can be partitioned into $K$ clusters $C_{1},\ldots,C_{K}$ so that, for any $k\in[K]$ and any $i,j\in C_{k}$, we have $\rho\left(r^{(i)}_{t},r^{(j)}_{t}\right)\leq\Delta$.
-For completeness, we present the SGD convergence analysis for strongly convex functions.
-The following analysis is well-known, but, compared with the most standard analysis, it works directly with the loss functions instead of distances to the optimum, thus avoiding losing the condition number in the following bounds.
-Lemma A.1.
+For completeness, we present the analysis of SGD convergence for functions satisfying $\mu$-PL condition.
+Lemma 2.
 Let $f$ be $L$-smooth and $\mu$-strongly convex.
@@
 $\leq f(\mathbf{x}_{t})-\frac{\eta}{2}\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}+\frac{L\eta^{2}}{2}\sigma^{2},$
-where in the last inequality we used $\eta\leq\frac{1}{L}$.
-Using the fact that $f$ is $\mu$-strongly-convex, we have $\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}\geq 2\mu(f(\mathbf{x}_{t})-f(\mathbf{x}^{*}))$, giving
+where in the last inequality we used $\eta\leq 1/L$.
+Using the fact that $f$ satisfies $\mu$-PL condition, we have $\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}\geq 2\mu(f(\mathbf{x}_{t})-f(\mathbf{x}^{*}))$, giving
 $\mathbb{E}_{t}\left[f(\mathbf{x}_{t+1})\right]$ $\leq f(\mathbf{x}_{t})-\eta\mu(f(\mathbf{x}_{t})-f(\mathbf{x}^{*}))+\frac{L\eta^{2}}{2}\sigma^{2}.$
@@
 $\mathbb{E}\left[f(\mathbf{x}_{T})-f(\mathbf{x}^{*})\right]$ $\leq\left(1-\eta\mu\right)^{T}(f(\mathbf{x}_{0})-f(\mathbf{x}^{*}))+\frac{L\eta}{2\mu}\sigma^{2}.\qed$
-The above result shows how the objective improves with each round: namely, after $R$, the non-stochastic part of the loss improves by a factor which depends on $R$ exponentially.
+The above result demosntrates how the objective improves with each round: namely, after $R$ iterations, the non-stochastic part of the loss improves by a factor which depends on $R$ exponentially.
 However, each data drift and each reclustering might potentially hurt the objective.
 Our next result bounds their effect on the loss.
-Lemma A.2.
+Lemma 3.
 Let $N$ be the number of clients, and $t+1$ be a fixed data drift event.
@@
 For a client $i$, let $k$ and $\ell$ be such that $i\in C_{k}\cap\bar{C}_{\ell}$. We rewrite $f^{(i)}_{t}(\bar{\mathbf{c}}_{\ell})-f^{(i)}_{t}(\bar{\mathbf{c}}_{\ell}^{*})$ as
-$f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell})-f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell}^{*})=(f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell})-f^{(i)}_{t+1}(\mathbf{c}_{k}^{*}))+(f^{(i)}_{t+1}(\mathbf{c}_{k}^{*})-f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell}^{*}))$
+$f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell})-f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell}^{*})=\left(f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell})-f^{(i)}_{t+1}(\mathbf{c}_{k}^{*})\right)+\left(f^{(i)}_{t+1}(\mathbf{c}_{k}^{*})-f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell}^{*})\right)$
 and bound each term separately.
@@
 Let $\mathbf{x}_{j}$ be the model of client $j$ before reclustering, i.e. $\mathbf{x}_{j}=\mathbf{c}_{k^{\prime}}$ where $j\in C_{k^{\prime}}$. Then,
-(2) $f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell})=f^{(i)}_{t+1}\Bigl{(}\operatorname*{avg}_{j\in\bar{C}_{\ell}}\mathbf{x}_{j}\Bigr{)}\leq\operatorname*{avg}_{j\in\bar{C}_{\ell}}f^{(i)}_{t+1}(\mathbf{x}_{j}),$
+$f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell})=f^{(i)}_{t+1}\Bigl{(}\operatorname*{avg}_{j\in\bar{C}_{\ell}}\mathbf{x}_{j}\Bigr{)}\leq\operatorname*{avg}_{j\in\bar{C}_{\ell}}f^{(i)}_{t+1}(\mathbf{x}_{j}),$ (1)
 where the last inequality follows by convexity of $f^{(i)}_{t}$.
@@
 $\geq f^{(i)}_{t}(\mathbf{c}_{k}^{*})-\theta(\Delta+\delta),$
-where the last inequality holds since otherwise $f^{(i)}_{t}(\mathbf{c}_{k}^{*})>f^{(i)}_{t}(\mathbf{x}_{t}^{(i,*)})+\theta\Delta$, which is impossible since $\mathbf{c}_{k}^{*}$ is the minimizer of $\operatorname*{avg}_{j\in C_{k}}f^{(j)}_{t}$ and $\operatorname*{avg}_{j\in C_{k}}f^{(j)}_{t}(\mathbf{x}_{t}^{(i,*)})\leq f^{(i)}_{t}(\mathbf{x}_{t}^{(i,*)})+\theta\Delta$ by Assumptions D and 3.
+where the last inequality holds since otherwise $f^{(i)}_{t}(\mathbf{c}_{k}^{*})>f^{(i)}_{t}(\mathbf{x}_{t}^{(i,*)})+\theta\Delta$, which is impossible since $\mathbf{c}_{k}^{*}$ is the minimizer of $\operatorname*{avg}_{j\in C_{k}}f^{(j)}_{t}$ and $\operatorname*{avg}_{j\in C_{k}}f^{(j)}_{t}(\mathbf{x}_{t}^{(i,*)})\leq f^{(i)}_{t}(\mathbf{x}_{t}^{(i,*)})+\theta\Delta$ by Assumptions D.
 Finally, we have $|f^{(i)}_{t}(\mathbf{c}_{k}^{*})-f^{(i)}_{t+1}(\mathbf{c}_{k}^{*})|<\theta\delta$ by Assumptions D and E. Combining the bounds concludes the proof. ∎
-Combining these results leads to our main convergence result.
-Theorem A.3.
+Combining the above statements leads to our main convergence result.
+Theorem 4.
 Let $N$ be the number of clients, and let their objective functions satisfy Assumptions B-D.
@@
 Then, for $\eta\leq 1/L$, for any data drift event $t$ of Algorithm 4 we have
-$\frac{1}{N}\sum_{k\in[K]}\sum_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t+1})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)\leq(1-\eta\mu)^{TR}\left(f(\mathbf{x}_{0})-f(\mathbf{x}^{*})\right)+\frac{L\eta}{2\mu}\left(\frac{\sigma^{2}+8L\theta\Delta}{M/K}+3\theta(\Delta+\delta)(1-\eta\mu)^{R}\right)$
+$\frac{1}{N}\sum_{k\in[K]}\sum_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t+1})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$ $\leq(1-\eta\mu)^{TR}\left(f(\mathbf{x}_{0})-f(\mathbf{x}^{*})\right)$
+$\quad+\frac{L\eta}{2\mu}\left(\frac{\sigma^{2}+8L\theta\Delta}{M/K}+3\theta(\Delta+\delta)(1-\eta\mu)^{R}\right)$
 Proof.
@@
 Since we sample $M/K$ clients from a cluster, the total variance – including the stochastic variance – is at most $\frac{\sigma^{2}+8L\theta\Delta}{M/K}$.
-Next, by Lemma A.2, the total objective increases by at most $3\theta(\Delta+\delta)$ after reclustering.
+Next, by Lemma 3, the total objective increases by at most $3\theta(\Delta+\delta)$ after reclustering.
 Since during $R$ rounds the non-stochastic part of each objective decreases by a factor of $(1-\eta\mu)^{R}$, we have
 $\frac{1}{N}\sum_{k\in[K]}|C_{k}|\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t+1}(\mathbf{c}^{(k)}_{t+1})-f^{(i)}_{t+1}(\mathbf{c}^{(k,*)}_{t+1})\right)$
-$\leq\frac{1}{N}\sum_{k\in[K]}|C_{k}|\left(\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)+3\theta(\Delta+\delta)\right)(1-\eta\mu)^{R}+\sum_{i=0}^{R}\left(1-\eta\mu\right)^{i}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$
-$=(1-\eta\mu)^{R}3\theta(\Delta+\delta)+(1-\eta\mu)^{R}\frac{1}{N}\sum_{k\in[K]}|C_{k}|\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)+\sum_{i=0}^{R}\left(1-\eta\mu\right)^{i}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$
-By Lemma A.1, the last (stochastic) term accumulates over all data drift events as $\frac{L\eta}{2\mu}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$.
+$\leq\frac{(1-\eta\mu)^{R}}{N}\sum_{k\in[K]}|C_{k}|\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})+3\theta(\Delta+\delta)\right)+\sum_{i=0}^{R}\left(1-\eta\mu\right)^{i}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$
+$=(1-\eta\mu)^{R}3\theta(\Delta+\delta)+\sum_{i=0}^{R}\left(1-\eta\mu\right)^{i}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$
+$\qquad+(1-\eta\mu)^{R}\frac{1}{N}\sum_{k\in[K]}|C_{k}|\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$
+By Lemma 2, the last (stochastic) term accumulates over all data drift events as $\frac{L\eta}{2\mu}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$.
 By a similar reasoning, $3\theta(\Delta+\delta)$ accumulates as $\frac{L\eta}{2\mu}\cdot 3\theta(\Delta+\delta)(1-\eta\mu)^{R}$.
@@
 ∎
-Appendix B Experiment settings
+Appendix B Additional background
+In this section, we provide the background of clustered federated learning and quantify the impacts of data drift.
+B.1 Clustered federated learning
+FL deployment often involves hundreds to thousands of available clients participating in the training (Bonawitz et al., 2019).
+For instance, Huang et al.
+(2019) have developed mortality and stay time predictors for the eICU collaborative research database (Pollard et al., 2018) containing data of 208 hospitals and more than 200,000 patients.
+Although thousands of devices may be available in a given round, FL systems typically select only a subset for training.
+For example, Google sets 100 as the target number of clients per training round when improving keyboard search suggestions (Yang et al., 2018).
+This selection process ensures a reasonable training time and accounts for diminishing returns, where including more clients does not accelerate convergence (Xu et al., 2024).
+Data heterogeneity is a major challenge in FL, as variations in local data volume and distribution among participants can lead to slow convergence and reduced model accuracy (Yang et al., 2021; Wen et al., 2023).
+Clustering is an effective strategy to mitigate the impact of heterogeneity by grouping statistically similar clients.
+Clients within the same cluster collaboratively train one cluster model, which then serves their inference requests.
+Prior works demonstrate that clustering achieve high accuracy and fast model convergence (Sattler et al., 2021; Mansour et al., 2020; Ghosh et al., 2020; Liu et al., 2023; Jothimurugesan et al., 2023).
+Our experiments with the Fielding system prototype mainly use label distribution vectors as the client representation, as they handle both label and covariate shifts and are lightweight.
+Although previous studies have examined label distribution vectors as leverage to address heterogeneity (Zhang and Lv, 2021; Zhang et al., 2022; Lee and Seo, 2023; Lee et al., 2021), Fielding distinguishes itself from existing works.
+Federated learning via Logits Calibration (Zhang et al., 2022) and FedLC (Lee and Seo, 2023) are orthogonal and investigate building a better global model through client model aggregation and loss function improvements.
+FedLabCluster (Zhang and Lv, 2021) and Lee et al.
+(2021) do not explicitly discuss data drift handling.
+Moreover, we present a theoretical framework with a convergence guarantee under data drift and re-clustering.
+B.2 Data drift affects static clustering
+Drifts happen naturally when clients have access to non-stationary streaming data (e.g., cameras on vehicles and virtual keyboards on phones) (Koh et al., 2021; Bhardwaj et al., 2022; Chen et al., 2020; Nandi and Xhafa, 2022; Marfoq et al., 2023).
+For example, when training an FL model for keyboard suggestions, drifts occur when many clients simultaneously start searching for a recent event (widespread drift), a few clients pick up a niche hobby (concentrated intense drift), or a new term appears when a product is launched (new label added).
+Data drift reduces the effectiveness of clustered FL as it increases data heterogeneity within clusters.
+We demonstrate this problem with the Functional Map of the World (FMoW) dataset (Christie et al., 2018).
+Recall from Section 3 that our preprocessed FmoW dataset contains time-stamped satellite images labeled taken on or after January 1, 2015, and each unique UTM zone metadata value corresponds to one client (302 clients in total).
+A day’s worth of data becomes available every two training rounds, and clients maintain images they received over the last 100 rounds.
+As mentioned in Section 2, we use mean client distance as the intra-cluster heterogeneity.
+Note that instead of first averaging within each cluster and then finding the mean of those cluster averages, we use the overall mean across all clients.
+This helps us avoid biased results arising from imbalanced cluster sizes.
+Note that in the baseline case without any clustering, we consider all clients as members of a "global cluster".
+We cluster clients based on their label distribution vectors at the starting round and track the overall heterogeneity by measuring the mean client distance described above.
+Figure 1 shows that initially (e.g., round 1), clustering reduces the per-cluster heterogeneity compared with no clustering (i.e., putting all clients in a global set).
+However, as data distribution shifts over time, static clustering increases per-cluster heterogeneity and soon gets close to the no clustering case at round 322.
+Appendix C Implementation
+Figure 11: The system architecture and workflow of Fielding.
+Blue arrows represent communication between clients and the coordinator.
+Black arrows indicate the workflow among different components on the same machine.
+We implemented Fielding in Python on top of FedScale (Lai et al., 2022), a state-of-the-art open-source FL engine.
+We follow FedScale’s design of having one centralized coordinator and client-specific executors.
+Figure 11 illustrates the end-to-end workflow of the Fielding system prototype with label distribution vectors as client representations.
+At the beginning of each training round, 1 client executors register with the coordinator to participate and report their local data distribution optionally.
+The coordinator’s cluster manager maintains clients’ distribution records and 2 moves clients to the closest cluster when their reported distribution drifts.
+After moving drifted clients individually, the cluster manager measures center shift distances and either triggers global re-clustering when any cluster center shifts significantly or finalizes clusters membership.
+In the case of global clustering, we use $K$-means clustering and determine $K$ using the silhouette method (Rousseeuw, 1987).
+The initial model of each newly created cluster is set as the average of its clients’ previous cluster model (see Algorithm 2).
+When client clustering is done, the coordinator notifies the client selector to 3 select a subset of clients to contribute to each cluster and 4 communicate the latest model parameters to the selected clients.
+5 Upon receiving a set of parameters, client executors conduct the training process over local data and 6 upload the updated model parameters to the coordinator.
+Finally, 7 the model manager aggregates individual models and stores the new cluster models.
+This process happens iteratively until the clients’ mean test accuracy reaches a target value.
+Fielding also regularly creates checkpoints for the models, clients’ metadata, and cluster memberships for future fine-tuning and failure recovery.
+Fielding’s overheads are mainly determined by the number and size of distribution vectors we need to re-cluster.
+In our largest evaluation setting where we train with 5078 clients on a dataset with 100 labels, per-client adjustment takes 2.0 seconds and global re-clustering takes 15.6 seconds on average.
+Storing the latest reported distribution vector for each client consumes $5078\times 100\times 4$B $\approx 1.9$MB of memory.
+Appendix D Experimental settings
 We conduct experiments on public datasets Functional Map of the World (FMoW) (Christie et al., 2018), Cityscapes (Cordts et al., 2016), Waymo Open (Sun et al., 2020), and Open Images (Kuznetsova et al., 2020).
@@
 In Table 2, we specify the training parameters including the total number of training rounds, learning rate, batch size, number of local steps, and the total number of clients selected to contribute in each round.
-Table 2. Dataset configurations.
+Table 1: Dataset configurations.
 Dataset Number of Clients Rounds Between New Data Arrivals
@@
 Open Images 5078 50
-Table 3. Training parameters.
-Dataset Total Rounds Learning Rate Batch Size Number of Local Steps Clients Selected per Round
+Table 2: Training parameters.
+Dataset Total Rounds Learning Rate Batch Size Number of Local Steps Participants per Round
 FMoW 2000 0.05 20 20 50
@@
 Open Images 400 0.05 20 20 200
+The URL, version information, and license of datasets we used are as follows:
+Appendix E Representations comparison
+E.1 Gradients as representation
+Table 3: Heterogeneity of all clients and intra-cluster heterogeneity after clustering (lower is better) on FMoW at various rounds.
+The left value denotes the pairwise L1 distance of distribution vectors, while the right represents the pairwise embedding squared Euclidean distance.
+Early on, gradient-based clustering is less effective due to the model being unstable, whereas label-based clustering decreases heterogeneity consistently.
+Notably, smaller distribution distances generally align with smaller embedding distances, suggesting that optimizing the label-based clustering objective also generates good embedding-based clusters.
+Total Rounds Un-Clustered Gradient-Based Clustering Label-Based Clustering
+100 1.81 46.33 1.82(+0.6%) 46.30(-0.06%) 1.65(-8.8%) 43.02(-7.14%)
+200 1.80 42.75 1.82(+1.1%) 42.53(-0.51%) 1.64(-8.9%) 41.52(-2.88%)
+500 1.78 36.99 1.71(-3.9%) 35.82(-3.16%) 1.62(-9.0%) 35.55(-3.89%)
+1000 1.74 31.85 1.66(-4.6%) 30.70(-3.61%) 1.56(-10.3%) 29.97(-5.90%)
+1500 1.76 32.35 1.63(-7.4%) 30.68(-5.16%) 1.60(-9.1%) 30.76(-4.91%)
+An issue with gradient-based re-clustering is that the clustering effectiveness is sensitive to the model quality.
+Table 3 presents the resulting average pairwise distribution vector L1 distance and embedding squared Euclidean distance when we perform gradient-based clustering after training a global model for various numbers of rounds on FMoW.
+The numbers in parentheses indicate the change in average distance relative to that of the global set.
+A negative value in the parentheses indicates that the generated clusters are overall less heterogeneous than the global set.
+When we perform gradient-based clustering with the global model at round 100 and round 200, the resulting clusters are as heterogeneous as the global set.
+As we postpone clustering to later rounds, gradient-based clustering achieves an increasingly larger reduction in mean client distance.
+This trend indicates that gradient-based clustering doesn’t perform well in earlier rounds when the global model we use to collect gradients from all clients has not converged.
+As mentioned in Section 2, we propose using label distribution vectors as client representations when we deal with label and covariate shifts.
+Label distribution vectors are available on all clients and enable us to promptly detect any drift by checking for distribution changes.
+Compared to gradient-based clustering, label-based doesn’t demand parameter or gradient transmission, doesn’t introduce additional computation tasks, and is not sensitive to the timing of client clustering.
+In Table 3, label-based clustering provides consistent heterogeneity reduction across rounds.
+Furthermore, as a label distribution vector typically has tens or hundreds of coordinates compared to millions in a full gradient, label-based clustering incurs significantly lower computational overhead.
+In our experiment of training ResNet-18 on the FMoW dataset, clustering all 302 clients using our label-based solution takes only 0.05 seconds while the gradient-based solution FlexCFL takes 37.92 seconds (note that FlexCFL already accelerates this process through dimensionality reduction using truncated Singular Value Decomposition).
+Figure 12: Fielding achieves high accuracy across client representations on Cityscapes.
+Figure 13: Gradient-based clustering handles concept drift better.
+In terms of overhead, both embedding- and gradient-based clustering introduce client-side computation, which can be mitigated by deploying a smaller shared model for representation collection, eliminating extra model download time.
+We adopt this approach when we gather the time-to-accuracy results of training ResNet-18 on Cityscapes with various representations (shown in Figure 13).
+We first construct a small training set by randomly sampling 200 images from each class.
+We then train a ResNet-18 model on this dataset for 300 epochs and broadcasts it so that clients store it locally and use it for gradient and embedding generation.
+Note that gradient-based Fielding still takes longer to finish 200 training rounds than other variants in Figure 13 due to the longer computational time of back propagation and longer transmission time of gradient vectors.
+Handling concept drifts where $P(y|x)$ changes requires loss-based representations, making gradients more appropriate than label distribution vectors or embeddings.
+Figure 13 shows the time-to-accuracy results with synthesized concept drifts on Cityscapes.
+Here we make all data samples available throughout the training, but introduce drift events by randomly choosing 50% clients and having each chosen client randomly pick two labels and swap their samples (i.e., if a client picks label $A$ and $B$, then all samples previously labeled $A$ are now labeled $B$, and samples labeled $B$ now have label $A$).
+Gradient-based Fielding manages to retain test accuracy under such aggressive concept drift, while label and embedding-based Fielding have similar results as the baseline without any clustering.
+This result highlights that gradient has the potential of addressing concept drifts while label distribution and embedding don’t.
+E.2 Input embeddings as representations
+Since covariate shift is defined as input distribution ($P(x)$) changing while the feature-to-label mapping ($P(y|x)$) remains constant, input embedding distance should be a proxy for label distribution distance.
+Hence, minimizing the distance between input embeddings naturally minimizes the label distribution distance as well.
+Table 3 supports this intuition by showing a correlation between the label distribution distance and the embedding distance.
+This observation suggests that using label distribution vectors or input embeddings as representations should have comparable performance on labeled datasets.
+As shown in Figure 13, both label distribution vectors and embeddings enable Fielding to outperform the no-clustering baseline (the sudden accuracy drop of the embedding-based curve around 150 minutes is the result of global re-clustering).
+Input embeddings have the potential of capturing label and covariate shifts on unlabeled data, which label distribution vectors are incapable of.
+Appendix F Additional results
+Figure 14: Fielding performance with different global re-clustering threshold ($\tau$).
+F.1 Global clustering threshold ablation study
+A tunable parameter of our Fielding prototype is the global re-clustering threshold $\tau$.
+We present the results on FMoW with $\tau$ being 0 (i.e., global re-clustering by default), $\frac{\theta}{6}$, $\frac{\theta}{3}$, $\frac{\theta}{2}$, and $\frac{2\theta}{3}$ in Figure 14 (recall that $\theta$ is the average distance between cluster centers).
+$\tau$ equals 0 leads to sudden accuracy drops that match our observations in Figure 2(a).
+$\tau=\frac{\theta}{6}$ stabilizes the system, but the final accuracy is similar to that of $\tau=0$.
+The largest threshold $\tau=\frac{2\theta}{3}$ performs the worst, matching Figure 1 that migrating individual clients without adequate global re-clustering leads to more heterogeneous clusters.
+We choose $\tau=\frac{\theta}{3}$ for our prototype as it gives the highest final accuracy.
+However, we acknowledge that the optimal threshold might be workload-specific.
+To address this, we propose making $\tau$ learnable during training: in the early rounds, we explore various thresholds and select those yielding the best accuracy; we can also periodically re-learn the threshold in later rounds.
+This ensures good performance across different datasets.
+F.2 Alternative global clustering threshold using pairwise distance
+To demonstrate that Fielding prototype performs empirically well with the alternate re-clustering thresholds presented in Appendix A, we re-run the Cityscapes task with re-clustering mechanism matching Line 9-11 of Algorithm 4.
+In specific, we measure clients’ pairwise distance as the L1 distance between their distribution vectors, and tune $\Delta$ adaptively.
+We start with $\Delta=c=0.1$.
+After each data drift event, we check whether global re-clustering has been triggered consecutive by two drift events.
+If so, we update $\Delta$ to $2\times\Delta$; otherwise, we update $\Delta$ to $\min(c,\Delta-c)$.
+Figure 16 shows that the prototype of the Fielding system works well with both center shift distance-based and client pairwise distance-based global re-clustering triggering conditions.
+Figure 15: Fielding performance with different global re-clustering triggering conditions (center shift distance as in Algorithm 2 or client pairwise distance as in Appendix A).
+Figure 16: Fielding performance with different distance metrics.
+F.3 Supporting diverse distance metric
+To demonstrate that Fielding is not tied to specific distance metric, we run the Cityscapes task with label distribution vectors as client representations, and both L1 and Jensen–Shannon distance (Lin, 1991) as the distance metric.
+Figure 16 shows that Fielding is compatible with different distance metrics chosen for a given client representation.

### v3

@@
 Data drift typically falls into three categories (Huyen, 2022): (1) label shift, where the output categories frequencies change but the characteristics associated with each category remain stable (Lipton et al., 2018; Garg et al., 2020); (2) covariate shift, where the distribution of the input features changes but the feature-label relationship remains constant (Ganguly and Aggarwal, 2023; Mallick et al., 2022); (3) concept shift, where the underlying input-output relationship changes, rendering a model trained on past data inaccurate on new data (Casado et al., 2022; Jothimurugesan et al., 2023).
-As data drift accumulates, intra-cluster heterogeneity increases, and clusters can become as diverse as the full client set (Jothimurugesan et al., 2023; Moreno-Torres et al., 2012).
+As data drift accumulates, intra‐cluster heterogeneity increases, and clusters can become as diverse as the full client set (Jothimurugesan et al., 2023; Moreno-Torres et al., 2012).
 Effective drift handling must address two key challenges: adapting to varying drift magnitudes while maintaining practical efficiency, and supporting diverse drift types.
@@
 Each representation has trade-offs: Label and embedding vectors capture only marginal distributions, while gradient-based features require additional computation, are sensitive to training noise, and evolve with model updates (Table 3 shows that gradient-based clustering generates better clusters as training progresses and the model becomes more stable).
-Selecting or combining these representations enables more reliable detection in all types of drift (see Appendix E).
+Selecting or combining these representations enables more reliable detection in all types of drift (see Appendix˜E).
 In this work, we present Fielding, a CFL framework that handles multiple types of data drift with low overhead.
@@
 Per-client adjustment methods, such as IFCA (Ghosh et al., 2020) and FlexCFL (Duan et al., 2021), maintain a fixed number of clusters and reassign drifted clients individually based on gradients or model weights.
-As mentioned in Section 1, such assignments become unstable and lead to poor clustering when many clients drift simultaneously.
+As mentioned in Section˜1, such assignments become unstable and lead to poor clustering when many clients drift simultaneously.
 Figure 1 (“Move Individuals”) highlights this behavior.
@@
 2.3 Algorithmic framework
-In this section, we describe our algorithm – presented in Algorithm 1 – in detail.
+In this section, we describe our algorithm – presented in Algorithm˜1 – in detail.
 Initially, Fielding clusters the clients using the $k$-means algorithm.
@@
 We assume that the algorithm is executed for $T$ global iterations, and during each iteration, the label distribution of each client might change, for instance by adding or removing data points.
-The first step of Fielding is to handle data drift (Algorithm 2).
+The first step of Fielding is to handle data drift (Algorithm˜2).
 Our goal is, by the end of each iteration, to maintain a good model for each cluster, with respect to its clients’ data.
@@
 This condition avoids frequent reclustering, which might require excessive resources and can adversely change the clients’ losses.
-Importantly, we recluster all the clients, not just available clients, which, as we show in Appendix B, improves the accuracy.
+Importantly, we recluster all the clients, not just available clients, which, as we show in Appendix˜B, improves the accuracy.
 For the global reclustering, we choose the number of clusters that gives the highest silhouette score.
@@
 3 for every iteration $t=0,\ldots,T-1$ do
-4       Handle data drift using Algorithm 2
+4   Handle data drift using Algorithm˜2
 5   Let $C_{1},\ldots,C_{K}$ be the clusters
@@
 2.4 Convergence
-In Appendix A, we analyze the performance of simplified version of our framework by bounding the average of clients’ loss functions at every iteration.
+In Appendix˜A, we analyze the performance of simplified version of our framework by bounding the average of clients’ loss functions at every iteration.
 The major challenge in the analysis is bounding the adverse effects of data drifts and reclustering.
 Data drifts change client datasets, hence changing their local objective functions.
-On the other hand, while reclustering is useful in the long run, its immediate effect on the objective can potentially be negative (see Figure 2(b)) due to the mixing of models from different clusters.
+On the other hand, while reclustering is useful in the long run, its immediate effect on the objective can potentially be negative (see Figure˜2(b)) due to the mixing of models from different clusters.
 We make the following assumptions.
@@
 Let $\mathbf{c}^{(k,*)}_{t}$ be the minimizer for cluster $k$ at iteration $t$.
-Then, under A, for $\eta\leq 1/L$, for any iteration $t$ we have
-$\frac{1}{N}\sum_{k\in[K]}\sum_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t+1})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$ $\leq(1-\eta\mu)^{tR}\left(f_{0}(\mathbf{x}_{0})-f_{0}(\mathbf{x}^{*})\right)$
+Then, under ˜A, for $\eta\leq 1/L$, for any iteration $t$ we have
+[…]
 $\quad+\frac{L\eta}{2\mu}\left(\frac{\sigma^{2}+8L\theta\Delta}{M/K}+3\theta(\Delta+\delta)(1-\eta\mu)^{R}\right)$
@@
 3 Evaluation
-We evaluate Fielding prototype on four FL tasks with label distribution vectors as client representations (see Appendix E for results with other representations).
+We evaluate Fielding prototype on four FL tasks with label distribution vectors as client representations (see Appendix˜E for results with other representations).
 End-to-end results highlight that Fielding improves final test accuracy by 1.9%-5.9% and is more stable than prior CFL methods.
@@
 3.1 End-to-end training performance
-Figure 4 shows that Fielding improves average client test accuracy by 5.9%, 1.9%, and 3.4% in FMoW, Cityscapes, and Waymo Open, respectively.
-These improvements are significant, given that baseline test accuracies range from 45% to 85%, and are comparable to the accuracy gains prior CFL works achieve on static data.
-When considering the global model’s final test accuracy as the target accuracy, Fielding offers a 1.27$\times$, 1.16$\times$, and 2.23$\times$ training speed-up, respectively (We define the point at which Fielding’s accuracy consistently surpasses the target accuracy as the moment it achieves the target accuracy).
-On the highly heterogeneous and dynamic Open Images trace, Fielding boosts accuracy by 17.9-26.1% in the final 100 rounds.
-While Auxo performs comparably or slightly better than the global model baseline, the benefits are limited: up to 0.9% accuracy gain and up to 1.05$\times$ convergence speedup.
-These results fall short of those reported in the original paper on static datasets, suggesting a decline in clustering effectiveness.
-This degradation stems from re-clustering only the selected participants after each round, ignoring drifted but unselected clients.
-This approach also leads to sudden accuracy drops on the biased Waymo Open dataset, where over 80% samples are cars, due to sudden model shifts when divergent clients are selected.
-FlexCFL shows no improvement in accuracy or convergence across all real-world datasets, as it migrates clients individually without adjusting the number of clusters.
-As Figure 1 shows, this approach leads to cluster heterogeneity approaching that of the global client set.
-FedDrift Comparison.
-FedDrift (Jothimurugesan et al., 2023) is another clustered FL method designed to handle drift.
-However, it is impractical for large-scale settings, as it (1) requires every client to train every cluster model in each round and (2) assumes all clients remain online and available for training.
-Therefore, we construct a small-scale setting with the FMoW dataset and 50 clients (the number of clients we selected per round in our original experiment).
-Figure 10 shows that Fielding achieves 2.5% higher final accuracy and reaches FedDrift’s final accuracy 2.01$\times$ faster.
+[…]
+[…]
+[…]
 3.2 Compatibility with other optimizations
@@
 We use the Oort selection algorithm (Lai et al., 2021) and a distance-based algorithm prioritizing clients closer to the distribution center as client selection examples.
-As shown in Figure 5, Fielding improves the final average test accuracy by 6.0% and 4.5% and reach the target accuracy 2.62×2.62$\times$2.62 × and 1.17×1.17$\times$1.17 × faster.
+As shown in Figure 5, Fielding improves the final average test accuracy by 6.0% and 4.5% and reach the target accuracy $2.62\times$ and $1.17\times$ faster.
 Note that Oort selection actively incorporates clients otherwise filtered out due to long response time, leading to a longer average round time.
 Figure 6 shows that Fielding works well with aggregation algorithm FedYogi (Reddi et al., 2021) and q-FedAvg (Li et al., 2020c).
-Fielding improves the final accuracy by 5.9% and 4.9% while giving a 1.34×1.34$\times$1.34 × and 1.25×1.25$\times$1.25 × speedup respectively.
+Fielding improves the final accuracy by 5.9% and 4.9% while giving a $1.34\times$ and $1.25\times$ speedup respectively.
 3.3 Robustness and sensitivity analysis
@@
 FedDrift (Jothimurugesan et al., 2023) re-clusters all clients by broadcasting all cluster models every round to compute client local training loss, leading to significant costs.
-Appendix B discusses other prior CFL works without drift handling and the impact of drifts on static clusters.
+Appendix˜B discusses other prior CFL works without drift handling and the impact of drifts on static clusters.
 5 Conclusion
@@
 Appendix A Proof of convergence
-1 input: the number of clusters $K$, the number of data drift events $T$, the number of sampled clients per round $M$, the number of rounds per data drift event $R$, reclustering threshold $\Delta$
+1input: the number of clusters $K$, the number of data drift events $T$, the number of sampled clients per round $M$, the number of rounds per data drift event $R$, reclustering threshold $\Delta$
 2 Initialize cluster models $\mathbf{c}^{(1)}_{0},\ldots,\mathbf{c}^{(K)}_{0}$ with the same model $\mathbf{x}_{0}$
@@
 7      Let $\mathbf{x}_{i}$ be the model corresponding to the cluster containing client $i$
-8      Let $r^{(1)}_{t},r^{(2)}_{t},\ldots$ be the client representations
+8  Let $r^{(1)}_{t},r^{(2)}_{t}$$\ldots$ be the client representations
 9 if there exists a cluster $C$ such that $\rho(r^{(i)}_{t},r^{(j)}_{t})>\Delta$ for clients $i,j\in C$ then
@@
 To simplify the proof, in the algorithm, we make the following changes compared with the algorithm from the main body: 1) we use the $k$-center clustering instead of $k$-means clustering; 2) we perform reclustering when the intra-clsuter heterogeneity exceeds a certain threshold.
-We compare this threshold condition with the one from the main body in Section F.2.
-We present the algorithm in Algorithm 4.
+We compare this threshold condition with the one from the main body in Section˜F.2.
+We present the algorithm in Algorithm˜4.
 After every data drift event, we first accept data changes for each client.
@@
 where the last inequality follows by convexity of $f^{(i)}_{t}$.
-By F, since all clients $\bar{C}_{\ell}$ belong to the same cluster, their representations differ by at most $\Delta$, and hence by Assumption D their local objectives differ by at most $\theta\Delta$.
+By ˜F, since all clients $\bar{C}_{\ell}$ belong to the same cluster, their representations differ by at most $\Delta$, and hence by Assumption D their local objectives differ by at most $\theta\Delta$.
 Therefore,
@@
 $\sum_{\ell\in[K]}\sum_{i\in\bar{C}_{\ell}}f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell})\leq\sum_{k\in[K]}\sum_{i\in C_{k}}f^{(i)}_{t+1}(\mathbf{c}_{k})$
-Hence, the sum of the first terms in Equation 1 over all $i$ can be bounded as
+Hence, the sum of the first terms in Equation˜1 over all $i$ can be bounded as
 $\sum_{k\in[K]}\sum_{i\in C_{k}}(f^{(i)}_{t+1}(\mathbf{c}_{k})-f^{(i)}_{t+1}(\mathbf{c}_{k}^{*})+\theta\Delta)$ $\leq\sum_{k\in[K]}\sum_{i\in C_{k}}(f^{(i)}_{t}(\mathbf{c}_{k})-f^{(i)}_{t}(\mathbf{c}_{k}^{*})+\theta(\Delta+2\delta)),$
@@
 Let $\mathbf{x}_{t}^{(i,*)}$ be the minimizer of $f^{(i)}_{t}$ and $\mathbf{x}_{t+1}^{(i,*)}$ be the minimizer of $f^{(i)}_{t+1}$. Then, by Assumptions D and F we have
-$f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell}^{*})$ $\geq f^{(i)}_{t+1}(\mathbf{x}_{t+1}^{(i,*)})$ ($\mathbf{x}_{t+1}^{(i,*)}$ is the minimizer of $f^{(i)}_{t+1}$)
+[…]
 $\geq f^{(i)}_{t}(\mathbf{x}_{t+1}^{(i,*)})-\theta\delta$ (Assumptions D and E)
-$\geq f^{(i)}_{t}(\mathbf{x}_{t}^{(i,*)})-\theta\delta$ ($\mathbf{x}_{t}^{(i,*)}$ is the minimizer of $f^{(i)}_{t}$)
+[…]
 $\geq f^{(i)}_{t}(\mathbf{c}_{k}^{*})-\theta(\Delta+\delta),$
@@
 Let $\mathbf{c}^{(k,*)}_{t}$ be the minimizer for cluster $k$ at data drift event $t$.
-Then, for $\eta\leq 1/L$, for any data drift event $t$ of Algorithm 4 we have
-$\frac{1}{N}\sum_{k\in[K]}\sum_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t+1})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$ $\leq(1-\eta\mu)^{TR}\left(f(\mathbf{x}_{0})-f(\mathbf{x}^{*})\right)$
+Then, for $\eta\leq 1/L$, for any data drift event $t$ of Algorithm˜4 we have
+[…]
 $\quad+\frac{L\eta}{2\mu}\left(\frac{\sigma^{2}+8L\theta\Delta}{M/K}+3\theta(\Delta+\delta)(1-\eta\mu)^{R}\right)$
@@
 First, we express the objective in terms of objectives for individual clusters:
-$\frac{1}{N}\sum_{k\in[K]}\sum_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$ $=\frac{1}{N}\sum_{k\in[K]}|C_{k}|\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$
+[…]
 We then analyze the convergence in each cluster.
@@
 Since we sample $M/K$ clients from a cluster, the total variance – including the stochastic variance – is at most $\frac{\sigma^{2}+8L\theta\Delta}{M/K}$.
-Next, by Lemma 3, the total objective increases by at most $3\theta(\Delta+\delta)$ after reclustering.
+Next, by Lemma˜3, the total objective increases by at most $3\theta(\Delta+\delta)$ after reclustering.
 Since during $R$ rounds the non-stochastic part of each objective decreases by a factor of $(1-\eta\mu)^{R}$, we have
-$\frac{1}{N}\sum_{k\in[K]}|C_{k}|\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t+1}(\mathbf{c}^{(k)}_{t+1})-f^{(i)}_{t+1}(\mathbf{c}^{(k,*)}_{t+1})\right)$
+[…]
 $\leq\frac{(1-\eta\mu)^{R}}{N}\sum_{k\in[K]}|C_{k}|\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})+3\theta(\Delta+\delta)\right)+\sum_{i=0}^{R}\left(1-\eta\mu\right)^{i}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$
 $=(1-\eta\mu)^{R}3\theta(\Delta+\delta)+\sum_{i=0}^{R}\left(1-\eta\mu\right)^{i}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$
-$\qquad+(1-\eta\mu)^{R}\frac{1}{N}\sum_{k\in[K]}|C_{k}|\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$
-By Lemma 2, the last (stochastic) term accumulates over all data drift events as $\frac{L\eta}{2\mu}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$.
+[…]
+By Lemma˜2, the last (stochastic) term accumulates over all data drift events as $\frac{L\eta}{2\mu}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$.
 By a similar reasoning, $3\theta(\Delta+\delta)$ accumulates as $\frac{L\eta}{2\mu}\cdot 3\theta(\Delta+\delta)(1-\eta\mu)^{R}$.
@@
 We demonstrate this problem with the Functional Map of the World (FMoW) dataset (Christie et al., 2018).
-Recall from Section 3 that our preprocessed FmoW dataset contains time-stamped satellite images labeled taken on or after January 1, 2015, and each unique UTM zone metadata value corresponds to one client (302 clients in total).
+Recall from Section˜3 that our preprocessed FmoW dataset contains time-stamped satellite images labeled taken on or after January 1, 2015, and each unique UTM zone metadata value corresponds to one client (302 clients in total).
 A day’s worth of data becomes available every two training rounds, and clients maintain images they received over the last 100 rounds.
-As mentioned in Section 2, we use mean client distance as the intra-cluster heterogeneity.
+As mentioned in Section˜2, we use mean client distance as the intra-cluster heterogeneity.
 Note that instead of first averaging within each cluster and then finding the mean of those cluster averages, we use the overall mean across all clients.
@@
 In the case of global clustering, we use $K$-means clustering and determine $K$ using the silhouette method (Rousseeuw, 1987).
-The initial model of each newly created cluster is set as the average of its clients’ previous cluster model (see Algorithm 2).
+The initial model of each newly created cluster is set as the average of its clients’ previous cluster model (see Algorithm˜2).
 When client clustering is done, the coordinator notifies the client selector to 3 select a subset of clients to contribute to each cluster and 4 communicate the latest model parameters to the selected clients.
@@
 This trend indicates that gradient-based clustering doesn’t perform well in earlier rounds when the global model we use to collect gradients from all clients has not converged.
-As mentioned in Section 2, we propose using label distribution vectors as client representations when we deal with label and covariate shifts.
+As mentioned in Section˜2, we propose using label distribution vectors as client representations when we deal with label and covariate shifts.
 Label distribution vectors are available on all clients and enable us to promptly detect any drift by checking for distribution changes.
@@
 In terms of overhead, both embedding- and gradient-based clustering introduce client-side computation, which can be mitigated by deploying a smaller shared model for representation collection, eliminating extra model download time.
-We adopt this approach when we gather the time-to-accuracy results of training ResNet-18 on Cityscapes with various representations (shown in Figure 13).
+We adopt this approach when we gather the time-to-accuracy results of training ResNet-18 on Cityscapes with various representations (shown in Figure˜13).
 We first construct a small training set by randomly sampling 200 images from each class.
 We then train a ResNet-18 model on this dataset for 300 epochs and broadcasts it so that clients store it locally and use it for gradient and embedding generation.
-Note that gradient-based Fielding still takes longer to finish 200 training rounds than other variants in Figure 13 due to the longer computational time of back propagation and longer transmission time of gradient vectors.
+Note that gradient-based Fielding still takes longer to finish 200 training rounds than other variants in Figure˜13 due to the longer computational time of back propagation and longer transmission time of gradient vectors.
 Handling concept drifts where $P(y|x)$ changes requires loss-based representations, making gradients more appropriate than label distribution vectors or embeddings.
-Figure 13 shows the time-to-accuracy results with synthesized concept drifts on Cityscapes.
+Figure˜13 shows the time-to-accuracy results with synthesized concept drifts on Cityscapes.
 Here we make all data samples available throughout the training, but introduce drift events by randomly choosing 50% clients and having each chosen client randomly pick two labels and swap their samples (i.e., if a client picks label $A$ and $B$, then all samples previously labeled $A$ are now labeled $B$, and samples labeled $B$ now have label $A$).
@@
 This observation suggests that using label distribution vectors or input embeddings as representations should have comparable performance on labeled datasets.
-As shown in Figure 13, both label distribution vectors and embeddings enable Fielding to outperform the no-clustering baseline (the sudden accuracy drop of the embedding-based curve around 150 minutes is the result of global re-clustering).
+As shown in Figure˜13, both label distribution vectors and embeddings enable Fielding to outperform the no-clustering baseline (the sudden accuracy drop of the embedding-based curve around 150 minutes is the result of global re-clustering).
 Input embeddings have the potential of capturing label and covariate shifts on unlabeled data, which label distribution vectors are incapable of.
@@
 F.2 Alternative global clustering threshold using pairwise distance
-To demonstrate that Fielding prototype performs empirically well with the alternate re-clustering thresholds presented in Appendix A, we re-run the Cityscapes task with re-clustering mechanism matching Line 9-11 of Algorithm 4.
+To demonstrate that Fielding prototype performs empirically well with the alternate re-clustering thresholds presented in Appendix˜A, we re-run the Cityscapes task with re-clustering mechanism matching Line 9-11 of Algorithm˜4.
 In specific, we measure clients’ pairwise distance as the L1 distance between their distribution vectors, and tune $\Delta$ adaptively.
@@
 If so, we update $\Delta$ to $2\times\Delta$; otherwise, we update $\Delta$ to $\min(c,\Delta-c)$.
-Figure 16 shows that the prototype of the Fielding system works well with both center shift distance-based and client pairwise distance-based global re-clustering triggering conditions.
-Figure 15: Fielding performance with different global re-clustering triggering conditions (center shift distance as in Algorithm 2 or client pairwise distance as in Appendix A).
+Figure˜16 shows that the prototype of the Fielding system works well with both center shift distance-based and client pairwise distance-based global re-clustering triggering conditions.
+Figure 15: Fielding performance with different global re-clustering triggering conditions (center shift distance as in Algorithm˜2 or client pairwise distance as in Appendix˜A).
 Figure 16: Fielding performance with different distance metrics.
@@
 To demonstrate that Fielding is not tied to specific distance metric, we run the Cityscapes task with label distribution vectors as client representations, and both L1 and Jensen–Shannon distance (Lin, 1991) as the distance metric.
-Figure 16 shows that Fielding is compatible with different distance metrics chosen for a given client representation.
+Figure˜16 shows that Fielding is compatible with different distance metrics chosen for a given client representation.